### v1

# BioMoDiffuse: Physics-Guided Biomechanical Diffusion for Controllable and Authentic Human Motion Synthesis

Abstract
Human motion generation holds significant promise in fields such as animation, film production, and robotics. However, existing methods often fail to produce physically plausible movements that adhere to biomechanical principles. While recent autoregressive and diffusion models have improved visual quality, they frequently overlook essential biodynamic features, such as muscle activation patterns and joint coordination, leading to motions that either violate physical laws or lack controllability. This paper introduces BioMoDiffuse, a novel biomechanics-aware diffusion framework that addresses these limitations. It features three key innovations: (1) A lightweight biodynamic network that integrates muscle electromyography (EMG) signals and kinematic features with acceleration constraints, (2) A physics-guided diffusion process that incorporates real-time biomechanical verification via modified Euler-Lagrange equations, and (3) A decoupled control mechanism that allows independent regulation of motion speed and semantic context. We also propose a set of comprehensive evaluation protocols that combines traditional metrics (FID, R-precision, etc.) with new biomechanical criteria (smoothness, foot sliding, floating, etc.). Our approach bridges the gap between data-driven motion synthesis and biomechanical authenticity, establishing new benchmarks for physically accurate motion generation.
Figure 1: Our model, BioMoDiffuse, generates physically realistic motions based on textual input. By embedding Euler-Lagrange dynamics into the generation process, the model produces physical-optimized trajectories that inherently satisfy physical laws, achieving superior motion authenticity compared to current SOTA approaches. Specifically, BiMoDiffuse mitigates common kinematic artifacts including foot sliding, body interpenetration, and motion discontinuities while ensuring temporal smoothness.
1 Introduction
Human motion generation has wide-ranging applications in animation, film, and robotics. In recent years, there has been growing focus on generating human motion conditioned on language, sound, and motion trajectories [68, 61, 12, 49]. Among these, language-guided motion generation stands out due to its high practical value, complexity, and the extensive foundational research supporting the field. Recent advancements [15, 48, 66, 17, 22, 24, 50, 88] in autoregressive and diffusion models have significantly improved the synthesis of diverse motion sequences while maintaining semantic consistency.
While existing methods have made strides in enhancing the visual plausibility of motion sequences, studies in biomechanics and animation highlight significant shortcomings[54, 10] of them in capturing key aspects of human motion. The absence of explicit modeling for biomechanical and dynamic features prevents current models from learning essential information, such as acceleration patterns and muscle activation levels. Consequently, generated motions are often rigid and physically unrealistic, deviating significantly from natural human movement in rhythm and joint coordination. Additionally, current evaluation metrics fail to quantitatively assess physical plausibility and motion velocity, while existing methods struggle with precise control over movement amplitude and speed. These limitations considerably diminish the practical value of motion generation in professional applications that require biomechanical accuracy.
To address these challenges, we propose BioMoDiffuse, a biomechanical-aware motion generation framework that seamlessly integrates musculoskeletal dynamics with diffusion processes. Our key innovation lies in embedding explicit physical supervision into the diffusion pipeline, leveraging a physics-aware network based on modified Euler-Lagrange equations and physics guidance to enforce biomechanical plausibility. Specifically, BioMoDiffuse introduces real-time biomechanical verification during denoising and a decoupled motion control system for independent regulation of velocity and semantics. Additionally, since current evaluation protocols neglect biomechanical-related aspects, we establish a set of comprehensive evaluation metrics emphasizing physical realism, combining biomechanical validity metrics (e.g., motion smoothness, foot sliding rate) with traditional generation metrics (FID, Diversity). Experiments on HumanML3D [14] and KIT-ML [51] demonstrate significant improvements over existing methods.
The main contributions of this paper are as follows:
  • A biomechanics-guided diffusion framework that integrates EMG signals and kinematic features with acceleration-based physical supervision, ensuring compliance with rigid-body dynamics .
  • A physics-aware network architecture based on Euler-Lagrange principles, incorporating a novel motion control system for enhancing controllability.
  • A comprehensive evaluation framework that emphasizes physical realism, unifying biomechanical metrics and traditional metrics.
2 Related Work
Physical plausible motion processing. In motion generation, reinforcement learning (RL) has been utilized to fine-tune models for physical plausibility [72, 89, 67]. For example, ReinDiffuse [18] and PhysDiff [77] extend MDM [66] to improve motion quality. However, these RL-based approaches predominantly focus on refining foot-ground contact using opaque, black-box reward functions, often overlooking broader rigid-body dynamics and biomechanical principles. The lack of transparency in their loss design further constrains interpretability and limits precise control over kinematic behaviors. Alternatively, some methods incorporate auxiliary constraints during training [75, 29, 53, 34]; for instance, OmniControl [75] leverages spatial signals to guide generation. Nonetheless, these approaches tend to prioritize controllability at the expense of enforcing comprehensive physical laws, disproportionately emphasizing foot-centric adjustments rather than ensuring holistic motion plausibility.
Physical plausibility is also addressed in other motion-related tasks that demand biomechanically authentic synthesized motion sequences, such as human pose estimation from videos [32, 37, 33, 63, 60]. Recent advances try to incorporate physics-based modeling [78, 58, 87, 86, 59, 52, 35, 74] to enhance the quality of the estimated motion sequences. However, these methods still face challenges of suboptimal precision and excessive jitter. Although techniques like Gaussian filtering can suppress high-frequency noise, they often compromise precision. Consequently, several studies [76, 7, 25, 79, 69, 30, 65] have explored pose refinement techniques to reduce temporal instability. For instance, SmoothNet [79] employs a parametric approach to optimize the temporal dimension, substantially enhancing motion quality.
Diffusion models. Diffusion models have demonstrated remarkable success in image generation, with various methods leveraging language guidance to produce high-quality images [55, 57, 81]. Among these approaches, DDPM [20] has emerged as the most widely adopted sampler, achieving impressive results. Building on DDPM, DDIM [62] reduces the number of inference steps while enabling deterministic sampling by adjusting the variance. The introduction of Classifier Guidance (CG) [9] empowered diffusion models to handle conditional generation tasks effectively, and CFG [19] further streamlined this process by replacing CG, enhancing the efficiency of guidance integration during training. In this paper, we observe that incorporating additional domain knowledge parallels the mechanisms of CG and CFG conceptually, leading us to re-derive the diffusion process accordingly.
In recent years, many studies have tackled the Diffusion Inversion problem, which provides a method for mapping real images to latent space, with broad applications in image editing [46, 71, 36, 26, 85, 8]. However, Diffusion Inversion methods have rarely been applied to tasks outside of image generation. We observe that both motion refinement and image editing aim to preserve the macroscopic structure while modifying finer details. Consequently, we are the first to apply DDIM Inversion to the task of motion refinement.
Text-driven human motion generation. Early approaches to text-driven human motion generation conceptualized motion as a form of language, adopting autoregressive methods commonly used in language models [2, 1, 48, 3]. Recognizing the potential of large language models, subsequent studies improved motion embeddings to enable the joint learning of motion and text, yielding promising results [15, 16, 80, 11, 73, 38, 42].
MDM [66] pioneered the application of diffusion models to human motion generation, while MLD [5], inspired by VAE [31], introduced a diffusion process within the latent space. MotionDiffuse [82] established a larger model framework, using cross-attention to link motion and text. Building on MotionDiffuse, ReMoDiffuse [83] incorporated retrieval-based motion databases and supplementary knowledge, resulting in significant performance improvements. MoFusion [43] and StableMoFusion [22], drawing inspiration from Stable Diffusion [55], utilized the UNet [56] architecture without relying on external knowledge, achieving state-of-the-art results. Notably, experiments with StableMoFusion revealed that DiT [47] and RetNet [64] underperformed relative to UNet, validating our choice of UNet as the backbone for the diffusion framework.
3 Method
Current approaches fail to capture the essential characteristics of human motion due to the absence of biomechanical constraints and rigid-body dynamics modeling, limiting their ability to learn authentic movement patterns. To address this gap, we present a novel motion generation framework that explicitly incorporates biomechanical principles and rigid-body kinematics, which is capable of synthesizing authentic physically plausible human motions. Our methodological contributions are as follows: Sec. 3.1 establishes the theoretical foundation by formalizing the biomechanical constraints; Sec. 3.2 introduces a physics-aware network architecture derived from Euler-Lagrange equations, ensuring dynamic compliance; Sec. 3.3 presents our biomechanics-guided diffusion based on joint probability distribution analysis. The complete motion generation framework is detailed in Sec. 3.4. Furthermore, Sec. 3.5 demonstrates that our proposed method, equipped with diffusion inversion, can be effectively utilized for physics-aware motion refinement.
Figure 2: Overview of BioMoDiffuse. At each diffusion timestep $t$, our framework processes the motion sequence $\Theta^{t}$ alongside text conditions. The text is encoded using a frozen CLIP encoder and concatenated with $\Theta^{t}$ as input to the UNet backbone. The light brown panel (left) illustrates our physics-aware network, which consists of: (1) An Accel Net that implements augmented Euler-Lagrange equations to predict biomechanically valid accelerations, and (2) An Autoregressive Block that models temporal dependencies. The training objectives combine: $\mathcal{L}_{\text{phys}}$ (from Accel Net), $\mathcal{L}_{\text{diffusion}}$ (reconstruction), and $\mathcal{L}_{\text{accel}}$ (acceleration supervision against ground truth). Structural details of the orange (autoregressive block) and light blue (Accel Net) components are provided in the right panels.
3.1 Preliminaries
Text-guided human motion generation requires an input text to condition the generation of a motion sequence $\Theta\in\mathbb{R}^{T\times D}$, where $T$ denotes the sequence length and $D$ represents the dimension of the joint representation. This joint representation can be converted into a coordinate representation $\Phi\in\mathbb{R}^{T\times 3d}$, where $d$ is the number of joints. The Euler-Lagrange equation is given by:
$F=M\cdot\ddot{\Phi_{t}}+C,$ (1)
where $F$ represents the internal and external contact forces, $M$ is the inertia matrix, and $C$ captures the combined effect of the Coriolis force and gravity. The term $F$ can be decomposed into internal force $I$ and the external contact force $E$. According to biomechanical principles, the internal force $I$ is generated by muscle movements and can be expressed as:
$I=R\cdot A(u),$ (2)
where $R$ is the muscle rotation matrix, $A(u)$ is the muscle activation force, and $u$ is the muscle activation coefficient. Using this decomposition, the Euler-Lagrange equation for human motion can be rewritten as Eq. 3. Please refer to Appendix D for details.
$\ddot{\Phi_{t}}=M^{-1}(E+C+R\cdot A(u)).$ (3)
3.2 Physics-Aware Network
The Euler-Lagrange equation-based network provides a stronger prior than conventional neural networks, as it inherently captures acceleration dynamics while reducing problem complexity. A crucial design here is to directly embed the constraints within the network architecture itself. In a physical sense, the Euler-Lagrange equation serves as a rigorous constraint for human motion: by framing motion generation as $Motion=Gen(text)$, this equation ensures the learned function aligns closely with real-world biomechanics; From a deep learning perspective, Euler-Lagrange equation networks offer a more robust alternative to traditional non-parametric constraints for handling diverse human skeleton data in supervised settings. Additionally, the inherent properties of Euler-Lagrange equations in our problem context guarantee convergence to a stable and unique solution.
Following the MIA dataset representation, a learnable mapping exists between human poses and the activation coefficients of primary muscles $u\in\mathbb{R}^{8}$, defined as $u=u(\Phi_{:t})$. The muscle rotation matrix $R$ depends solely on pose rotation, while the inertia matrix $M$—a symmetric matrix—is intrinsic to the body and determined by position. The Coriolis force and gravity are formulated as:
$C=\dot{M}(\Phi_{t})\dot{\Phi_{t}}-\frac{1}{2}\dot{\Phi_{t}}^{T}\frac{\partial M}{\partial\Phi_{t}}\dot{\Phi_{t}}+G(\Phi_{t}),$ (4)
which are also pose-dependent. External contact forces $F$ (e.g., elastic/resistive forces) depend on both position and velocity. Thus, Eq. 3 is reformulated as:
$\ddot{\Phi}_{t}=M^{-1}(\Phi_{:t})\left(F(\Phi_{:t})+C(\Phi_{:t})+R(\Phi_{:t})\odot A(u)\right),$ (5)
where $\odot$ denotes element-wise multiplication. We implement three neural networks to approximate $M$, $F$, and $C$, with $M$ constrained as symmetric. The term $R\odot A(u)$ is modeled via cross-attention, enabling lightweight networks to predict pose acceleration. Finally, leveraging autoregressive processing, the predicted $\ddot{\Phi}_{t}$ and speed control signal are integrated as inputs to derive the optimized $\Theta_{t}$, ensuring temporal coherence and biomechanical fidelity.
3.3 Physics Guidance
We introduce additional guidance into the diffusion process to ensure that acceleration remains physically meaningful and achieve more stable training and faster convergence. Our network explicitly treats the acceleration $\ddot{\Phi_{t}}$ as an intermediate variable. We aim for the acceleration $\ddot{\Phi_{t}}$ in our physics-based network to converge to the real result as quickly as possible. To achieve this, we impose a supervised constraint on the acceleration during the diffusion process. Specifically, at each reverse diffusion step, we consider extending the distribution at step $t$, denoted by $\Theta^{t}$, into a joint distribution $S^{t}=\{\Theta^{t},\ddot{\Phi}^{t}\}$. According to DDPM theory, when performing diffusion on $S^{t}$, the model’s loss function can be written as:
$\min E_{Q(S^{0}:S^{T})}\left[-\log P_{\theta}(S^{0}|S^{1})\right.$ (6)
$+\sum_{t=2}^{T}\text{KL}(Q(S^{t-1}|S^{t},S^{0})\|\left.P_{\theta}(S^{t-1}|S^{t}))\right].$
We argue that, during the denoising process, the quantities $\Theta^{t-1},\ddot{\Phi}^{t-1}$ at step $t-1$ are unknown and do not satisfy the interpolation relationship. However, the known quantities $\Theta^{t},\ddot{\Phi}^{t}$ at step $t$ do satisfy this relationship. This leads to the following:
$\log P_{\theta}(S^{t-1}|S^{t})=\log P_{\theta}(\Theta^{t-1},\ddot{\Phi}^{t-1}|\Theta^{t},\ddot{\Phi}^{t})$ (7)
$=\log P_{\theta}(\Theta^{t-1},\ddot{\Phi}^{t-1}|\Theta^{t})$
$=\log P_{\theta}(\Theta^{t-1}|\ddot{\Phi}^{t-1},\Theta^{t})+\log P_{\theta}(\ddot{\Phi}^{t-1}|\Theta^{t}).$
In the expression in Eq. 7, $\log P_{\theta}(\Theta^{t-1}|\ddot{\Phi}^{t-1},\Theta^{t})$ resembles the condition $c$ in CG and CFG, while $\log P_{\theta}(\ddot{\Phi}^{t-1}|\Theta^{t})$ can be obtained using the physical network we propose. Thus, during training, we omit the physical network with probability $P_{c}$, and use it with probability $1-P_{c}$. We set $\ddot{\Phi}^{t-1}=\text{PhysNet}(\Theta^{t})$, and the predicted acceleration is compared to the ground truth acceleration, introducing a physics loss $L_{\text{phys}}=\text{MSE}(\text{PhysNet}(\Theta^{t}),\ddot{\Phi_{\text{gt}}}^{t})$. To further enhance acceleration accuracy, we design the following additional loss term for supervision:
$L_{\text{accel}}=\text{MSE}(\ddot{\Phi}^{t},\ddot{\Phi_{\text{gt}}}^{t}).$ (8)
Thus, the total loss function during training can be written as:
$L=L_{\text{diffusion}}+L_{\text{accel}}+L_{\text{phys}}.$ (9)
3.4 Model Architecture
Table 1: Quantitative evaluation on the HumanML3D and KIT-ML test set. $\pm$±indicates a 95% confidence interval. Bold face indicates the best result, while underscore refers to the second best. We reimplemented StableMoFusion based on the official code.
[…]
Top 1 Top 2 Top 3
KIT- ML MDM $0.164^{\pm.004}$ $0.291^{\pm.004}$ $0.396^{\pm.004}$ $0.497^{\pm.021}$ $9.191^{\pm.022}$ $10.85^{\pm.109}$ $1.91^{\pm.214}$
MLD $0.390^{\pm.008}$ $0.609^{\pm.008}$ $0.734^{\pm.027}$ $0.404^{\pm.027}$ $3.204^{\pm.027}$ $10.80^{\pm.117}$ $2.19^{\pm.071}$
T2M $0.370^{\pm.005}$ $0.569^{\pm.007}$ $0.693^{\pm.007}$ $2.770^{\pm.109}$ $3.401^{\pm.008}$ $10.91^{\pm.119}$ $1.48^{\pm.065}$
MotionDiffuse $0.417^{\pm.004}$ $0.621^{\pm.004}$ $0.739^{\pm.004}$ $1.954^{\pm.062}$ $2.958^{\pm.005}$ $11.10^{\pm.143}$ $0.73^{\pm.013}$
T2M-GPT $0.416^{\pm.006}$ $0.627^{\pm.006}$ $0.745^{\pm.006}$ $0.514^{\pm.029}$ $3.007^{\pm.023}$ $10.92^{\pm.108}$ $1.57^{\pm.039}$
ReMoDiffuse $0.427^{\pm.014}$ $0.641^{\pm.004}$ $0.765^{\pm.055}$ $0.155^{\pm.006}$ $2.814^{\pm.012}$ $10.80^{\pm.105}$ $1.24^{\pm.028}$
StableMoFusion $0.445^{\pm.006}$ $0.660^{\pm.005}$ $0.782^{\pm.004}$ $0.258^{\pm.029}$ $2.832^{\pm.014}$ $10.94^{\pm.077}$ $1.36^{\pm.062}$
Ours $0.448^{\pm.008}$ $0.666^{\pm.005}$ $0.788^{\pm.005}$ $0.211^{\pm.101}$ $2.772^{\pm.017}$ $11.11^{\pm.094}$ $1.38^{\pm.050}$
Human ML3D MDM $0.320^{\pm.005}$ $0.498^{\pm.004}$ $0.611^{\pm.007}$ $0.544^{\pm.044}$ $5.566^{\pm.027}$ $9.559^{\pm.086}$ $2.799^{\pm.072}$
MLD $0.481^{\pm.003}$ $0.673^{\pm.003}$ $0.772^{\pm.002}$ $0.473^{\pm.013}$ $3.196^{\pm.010}$ $9.724^{\pm.082}$ $2.413^{\pm.079}$
T2M $0.457^{\pm.002}$ $0.639^{\pm.003}$ $0.743^{\pm.003}$ $1.067^{\pm.002}$ $3.340^{\pm.008}$ $9.188^{\pm.002}$ $2.090^{\pm.083}$
MotionDiffuse $0.491^{\pm.001}$ $0.681^{\pm.001}$ $0.782^{\pm.001}$ $0.630^{\pm.001}$ $3.113^{\pm.001}$ $9.410^{\pm.049}$ $1.553^{\pm.042}$
T2M-GPT $0.491^{\pm.003}$ $0.680^{\pm.003}$ $0.775^{\pm.002}$ $0.116^{\pm.004}$ $3.118^{\pm.011}$ $9.761^{\pm.081}$ $1.856^{\pm.011}$
ReMoDiffuse $0.510^{\pm.005}$ $0.698^{\pm.006}$ $0.795^{\pm.004}$ $0.103^{\pm.004}$ $2.974^{\pm.016}$ $9.018^{\pm.075}$ $1.795^{\pm.043}$
StableMoFusion $0.545^{\pm.003}$ $0.743^{\pm.003}$ $0.833^{\pm.002}$ $0.133^{\pm.005}$ $2.814^{\pm.008}$ $9.704^{\pm.065}$ $1.844^{\pm.060}$
Ours $0.547^{\pm.003}$ $0.743^{\pm.002}$ $0.835^{\pm.002}$ $0.071^{\pm.003}$ $2.784^{\pm.008}$ $9.567^{\pm.086}$ $1.919^{\pm.063}$
Our biomechanics-aware diffusion framework is built upon three core components, as shown in Fig. 2. The architecture seamlessly integrates biomechanical priors with diffusion dynamics through:
Biomechanical Encoding. Using the MIA dataset [6], we pre-train a neuromuscular encoder that establishes a dynamical mapping $\mathcal{F}:\Theta\rightarrow u$ between joint angles $\Theta$ and muscle activation coefficients $u$. This encoder remains differentiable throughout end-to-end training.
Physics-Conditioned Diffusion. The UNet-based diffusion backbone processes motion sequences $\mathbf{x}_{t}$ at timestep $t$ through:
$z_{\theta}=\text{BioMoDiffuse}\big{(}\text{CLIP}(\mathbf{c}),\mathbf{x}_{t},\mathcal{F}(\Theta)\big{)},$ (10)
where text conditioning $\mathbf{c}$ is incorporated via cross-attention layers. We adopt a two-phase training process: initial diffusion followed by physics-constrained fine-tuning (which introduces the physics loss $\mathcal{L}_{\text{phys}}$).
Controllable Generation. Our decoupled control mechanism enables independent manipulation of motion dynamics and semantics through: (1) classifier-free guidance and physics guidance weights, and (2) the injection of biomechanical constraints via the pre-trained encoder.
3.5 Pose Refinement with Diffusion Inversion
Figure 3: Refining pose using our method. Building on existing pose estimation models, we integrate BioMoDiffuse with DDIM inversion [62]. By adopting an inversion-then-forward pipeline, our approach significantly improves the smoothness of generated SMPL sequences while enhancing the fitting accuracy of pose estimation.
Our proposed framework can be readily adapted to refine pose estimates from existing pose estimation methods. The key insight is that our physics-guided diffusion model is capable of generating physically plausible motions starting from latent noise. Given the initial estimates of human poses, we first employ diffusion inversion—a method that reverses the usual diffusion process—to recover the latent noise corresponding to the poses. We then feed this noise into the physics-guided diffusion model, which injects biomechanical constraints to produce the refined poses that are more authentic compared to original estimates. Specifically, given the initial SMPL [39] parameters $\mathbf{x}_{0}$ from pose estimation, we first invert them to timestep $t$ using DDIM sampling:
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\epsilon_{\theta}(\mathbf{x}_{0},t),$ (11)
where $t$ controls the balance between information preservation and noise introduction. The inverted latent $\mathbf{x}_{t}$ then undergoes forward diffusion with our physics guidance to produce the refined poses:
$\hat{\mathbf{x}_{0}}=f_{\theta}(\mathbf{x}_{t},t).$ (12)
Crucially, our parametric implementation inherits the physics constraints from Secs. 3.2 and 3.3 without the need for additional tuning, demonstrating a unified biomechanical framework across both motion generation and refinement tasks.
4 Experiments
4.1 Datasets and Metrics
We conducted training and testing on our model using both the KIT-ML [51] and HumanML3D [14] datasets. The KIT-ML dataset contains 3,911 motions and 6,363 natural language annotations, while HumanML3D is derived from and reannotates the HumanAct12 [13] and AMASS [44] datasets, comprising a total of 14,616 motion sequences and 44,970 descriptions. We follow the evaluation metrics used in previous motion generation works [66, 82, 22], including Frechet Inception Distance (FID), Precision, Diversity, Multimodality, and Multi-Modal Distance. Additionally, we consider physics-related metrics such as skating ratio, ground floating, ground penetration, and foot clipping, as used in prior studies [18, 75, 77]. Furthermore, we introduce two novel metrics based on Fourier analysis: Velocity Consistency (VC) and Acceleration Consistency (AC). Our motivation stems from the observation that while the Smoothness metric captures global temporal coherence, it overlooks the inherent kinematic relationship between stable body parts and agile extremities in human motion. For a motion sequence represented as $\Phi\in\mathbb{R}^{T\times 3d}$, the metrics are formally defined as:
VC $=\frac{1-\|\text{FFT}(\dot{\Phi})_{[:thre_{\text{low}}]}\|-\|\text{FFT}(\dot{\Phi})_{[thre_{\text{high}}:]}\|}{\|\text{FFT}(\dot{\Phi})_{[:thre_{\text{low}}]}\|+\|\text{FFT}(\dot{\Phi})_{[thre_{\text{high}}:]}\|},$ (13)
AC $=\frac{1-\|\text{FFT}(\ddot{\Phi})_{[:thre_{\text{low}}]}\|-\|\text{FFT}(\ddot{\Phi})_{[thre_{\text{high}}:]}\|}{\|\text{FFT}(\ddot{\Phi})_{[:thre_{\text{low}}]}\|+\|\text{FFT}(\ddot{\Phi})_{[thre_{\text{high}}:]}\|}.$
Spectral ratio analysis enables VC and AC to effectively characterize overall motion patterns by precisely quantifying the balance between high-frequency components (associated with rapid limb movements) and low-frequency components (correlated with stable torso dynamics). This frequency-domain approach provides a more accurate assessment of motion physicality compared to traditional smoothness metrics. Since the dataset includes motions like “climbing stairs” (a foot is floating from the ground) and “jumping while spinning” (abrupt acceleration), all these physics-based metrics should be made as consistent as possible with ground-truth instead of being blindly minimized to meet the characteristics of dataset.
4.2 Quantitative Results
Figure 4: User Study Results.
User Study.
We asked participants to rate four aspects of the generated motions on a 1–5 scale: “Generation Quality,” “Motion Smoothness,” “Physical Plausibility,” and “Text Alignment.” Using the HumanML3D test set. Thirty users were invited to compare the motion generation results of our model, StableMoFusion, and Ground Truth across 20 prompts. The results demonstrate that our model exhibits superior performance in terms of generation quality, motion smoothness, physical plausibility, and semantic alignment compared to StableMoFusion. Furthermore, on certain examples, our method achieves results that are comparable to the ground truth, showcasing its competitive performance.
Table 2: Quantitative evaluation of physical plausible metrics on the HumanML3D and KIT-ML test set. All experiments were repeated 20 times with different random seeds, and the mean results are reported in the tables. The results for T2M, and MotionDiffuse on HumanML3D is sourced from ReinDiffuse [18].
[…]
KIT- ML Real Motions 0.352 1.379 0.926 3.06 0.550 0.309 0.208 0.001
T2M $0.270$ $1.519$ $0.748$ $4.90$ $0.442$ $0.324$ $0.180$ $\bm{0.001}$
MotionDiffuse $0.258$ $1.861$ $1.177$ $1.92$ $0.136$ $0.808$ $0.490$ $0.005$
T2M-GPT $0.238$ $\bm{1.420}$ $0.832$ $3.97$ $0.503$ $0.267$ $0.188$ $0.002$
ReMoDiffuse $2.826$ $7.993$ $1.415$ $0.29$ $0.254$ $0.789$ $\bm{0.199}$ $0.006$
StableMoFusion $0.684$ $2.496$ $0.986$ $1.80$ $0.496$ $0.315$ $0.255$ $0.002$
Ours $\bm{0.426}$ $1.736$ $\bm{0.965}$ $\bm{2.25}$ $\bm{0.522}$ $\bm{0.306}$ $0.237$ $0.003$
Human ML3D Real Motions 0.124 0.682 0.358 2.60 0.205 0.057 0.000 0.000
T2M $0.326$ $2.826$ $0.301$ $1.13$ $0.110$ $0.217$ $0.235$ $0.263$
MotionDiffuse $0.082$ $0.963$ $0.332$ $1.17$ $0.031$ $0.426$ $0.386$ $0.217$
T2M-GPT $0.178$ $1.459$ $\bm{0.354}$ $1.28$ $0.192$ $0.052$ $0.022$ $\bm{0.001}$
ReMoDiffuse $2.353$ $7.163$ $0.605$ $0.15$ $0.040$ $0.376$ $0.019$ $0.008$
StableMoFusion $0.292$ $2.220$ $0.448$ $0.74$ $0.234$ $0.074$ $0.043$ $0.002$
Ours $\bm{0.135}$ $\bm{0.959}$ $0.385$ $\bm{1.45}$ $\bm{0.204}$ $\bm{0.056}$ $\bm{0.013}$ $0.002$
Main Results.
Tab. 1 and Tab. 2 compare our method with existing approaches across both physical plausibility metrics (e.g., smoothness, foot sliding) and traditional generation metrics (FID, Diversity). Our model achieves state-of-the-art results in biomechanical authenticity while simultaneously advancing traditional motion quality, outperforming prior works across multiple dimensions. Experiments reveal critical limitations in existing paradigms: GPT-style generative models tend to over-smooth motions, failing to capture realistic dynamics, while diffusion-based methods exhibit temporal discontinuity—likely due to their implicit treatment of spatiotemporal relationships without explicit temporal modeling. By integrating our physics-aware network into the diffusion framework, we address these issues: the network’s frame-wise acceleration recalculation enhances temporal dependency learning, directly improving physical metrics. Concurrently, the introduction of biomechanical priors reduces solution space complexity, enabling superior performance in traditional metrics without compromising motion diversity.
Figure 5: Qualitative comparison of the state-of-the-art methods. Quantitative evaluations demonstrate that BioMoDiffuse achieves superior language-motion alignment, enhanced temporal smoothness, and improved physical plausibility compared to current approaches.
4.3 Ablation Study
Table 3: Ablation Study on traditional metrics. we evaluate the impact of adding $L_{accel}$ loss, incorporating the Physical Network, and introducing Physics Guidance on the performance of our method. All experiments were repeated 20 times with different random seeds, and the mean results are reported in the tables.
[…]
Top 1 Top 2 Top 3
Ours $\times$ $\times$ $\times$ $0.090$ $2.824$ $9.590$ $1.849$ $0.542$ $0.739$ $0.832$
$\checkmark$ $\times$ $\times$ $0.129$ $2.835$ $\bm{9.659}$ $1.882$ $0.537$ $0.734$ $0.827$
$\checkmark$ $\checkmark$ $\times$ $\bm{0.066}$ $\bm{2.783}$ $9.596$ $\bm{1.948}$ $\bm{0.549}$ $\bm{0.746}$ $\bm{0.836}$
$\checkmark$ $\checkmark$ $\checkmark$ $0.071$ $2.784$ $9.567$ $1.919$ $0.547$ $0.743$ $0.835$
Table 4: Ablation Study on physical plausible metrics.
[…]
GT - - - $0.124$ $0.682$ $0.358$ $2.60$ $0.205$ $0.057$ $0.000$ $0.000$
Ours $\times$ $\times$ $\times$ $0.292$ $2.220$ $0.448$ $0.74$ $0.234$ $0.074$ $0.043$ $0.002$
$\checkmark$ $\times$ $\times$ $0.144$ $1.022$ $0.437$ $1.24$ $0.228$ $0.076$ $0.052$ $0.003$
$\checkmark$ $\checkmark$ $\times$ $0.135$ $\bm{0.954}$ $0.421$ $1.39$ $0.227$ $0.060$ $0.026$ $0.002$
$\checkmark$ $\checkmark$ $\checkmark$ $\bm{0.135}$ $0.959$ $\bm{0.385}$ $\bm{1.45}$ $\bm{0.204}$ $\bm{0.056}$ $\bm{0.013}$ $\bm{0.002}$
Our method introduces three key innovations to the diffusion framework: acceleration loss, a physics network grounded in Euler-Lagrange equations, and physics-guided sampling. Ablation studies under identical training settings Tab. 3 and Tab. 4 validate their contributions. While the acceleration loss enhances kinematic feature extraction by expanding the motion representation dimension from 263 to 329, it initially degrades generation quality (e.g., FID deteriorates by 43.3%) due to increased solution space complexity. The proposed physics network addresses this by embedding biomechanical constraints, reformulating motion generation as a constrained optimization problem. This reduces spatial complexity while enforcing rigid-body dynamics, leading to a 48.8% FID improvement with only a 6.68% parameter increase over the UNet baseline. Crucially, the network’s Euler-Lagrange foundation enables simultaneous gains in physical plausibility and generation quality. Finally, physics-guided sampling prioritizes biomechanical fidelity during denoising, stabilizing acceleration fitting and further improving physical metrics, albeit with a slight trade-off in generation scores—a phenomenon consistent with CFG principles.
4.4 Motion Refining Results
Table 5: Comparative results for pose refinement. We trained and tested the WHAM [60] baseline and WHAM $+$ our method using the official code on the 3DPW [70] and RICH [21] datasets. Bold numbers denote the most accurate method in each column. Accel is in $m/s^{2}$, all other errors are in $mm$.
[…]
3DPW WHAM $36.34$ $61.46$ $70.64$ $6.52$
+ Inversion $\bm{36.13}$ $\bm{61.11}$ $\bm{70.52}$ $\bm{5.55}$
RICH WHAM $50.67$ $83.00$ $95.32$ $5.35$
+ Inversion $\bm{50.60}$ $\bm{81.80}$ $\bm{94.84}$ $\bm{4.06}$
As shown in Tab. 5, we validate our model’s motion refinement capability using WHAM [60], a state-of-the-art pose estimation framework that leverages temporal modeling to achieve leading results on metrics like acceleration error. By integrating our physics-aware network with WHAM via diffusion inversion, we significantly improve motion smoothness while also advancing pose prediction accuracy to new state-of-the-art levels. This indicates that our model, equipped with diffusion inversion, can be applied after any pose estimation module to refine the motion sequence, resulting in more fluent and physically plausible results. These results demonstrate that our physical network effectively enhances temporal continuity and overall robustness, even without acceleration loss supervision. This dual capability underscores the versatility of our approach, showcasing its ability to refine both generated and estimated motions by enforcing physical plausibility at the intersection of kinematics and dynamics.
4.5 Ablation Experiments on Speed Control Signal
Table 6: Ablation Study on weight of Speed Control Signal. We evaluate the impact of different Speed Control Signal on generation quality and physical plausibility metrics. All experiments use the HumanML3D dataset.
[…]
GT $0.002$ $2.974$ $0.797$ $0.124$ $0.682$ $0.358$ $2.60$ $0.205$ $0.057$ $0.000$ $0.000$
$0.5$ $0.479$ $2.860$ $0.818$ $0.148$ $1.085$ $0.294$ $1.83$ $0.096$ $0.070$ $0.012$ $0.001$
$0.8$ $0.114$ $2.782$ $0.834$ $0.138$ $0.989$ $0.356$ $1.56$ $0.165$ $0.061$ $0.013$ $0.002$
$1.0$ $0.071$ $2.784$ $0.835$ $0.135$ $0.959$ $0.385$ $1.45$ $0.204$ $0.056$ $0.013$ $0.002$
$1.5$ $0.085$ $2.824$ $0.830$ $0.134$ $0.937$ $0.428$ $1.32$ $0.276$ $0.050$ $0.013$ $0.002$
$2.0$ $0.124$ $2.869$ $0.824$ $0.136$ $0.942$ $0.447$ $1.26$ $0.321$ $0.045$ $0.013$ $0.002$
Our physical network incorporates a speed control signal within the autoregressive block. Experiments reveal that applying this signal only during inference (via input weighting) dynamically scales motion amplitude, with stronger signals producing larger movements. However, three critical observations emerge: (1) Traditional generation metrics (e.g., FID) remain unaffected by signal strength, aligning with the convention that model parameters require consistency between training and testing; (2) Certain physical metrics, such as ground penetration, foot clipping, and velocity/acceleration consistency, also show minimal sensitivity; (3) Metrics tied to acceleration magnitude—smoothness and skating ratio—decrease with stronger signals, while ground floating increases. We attribute this to the explicit acceleration modeling in our framework: since the network learns absolute acceleration values (rather than relative relationships), scaling the input signal alters acceleration magnitudes, thereby directly impacting physics-based metrics. This demonstrates our model’s unique ability to modulate motion amplitude through acceleration control while preserving generation quality—a trade-off avoided by conventional methods. Furthermore, these findings guide targeted optimization of physical plausibility by tuning acceleration-related parameters.
5 Conclusion
In this paper, we present BioMoDiffuse, a motion generation model built upon a physical network derived from extended Euler-Lagrange equations. This model incorporates the physical network into pose estimation using DDIM Inversion. BioMoDiffuse combines physics guidance with a state-of-the-art UNet backbone architecture, while simultaneously learning from electromyographic (EMG) signals and acceleration data. Experimental results indicate that BioMoDiffuse surpasses previous models in both generation quality and physical plausibility. However, due to current dataset limitations, BioMoDiffuse has not yet been tested on higher-frame-rate datasets with more comprehensive annotations. We leave it for future exploration.
Supplementary Material
Video:
We have provided supplemental videos in our supplemental material. In the supplementary video, we demonstrate the following: 1) the differences between the human skeletons generated by our method and those by StableMoFusion [22], 2) the differences in the skinned human models generated by our method and StableMoFusion, and 3) the text prompts used to generate these motions. All generations were produced using the same code under identical random seed conditions (seed=0). We reconstructed the meshes following the rendering approach provided by MLD [5].
Code:
The code has also been included in the supplementary materials, and detailed execution instructions can be found in the accompanying README file.
Appendix A Evaluation Metrics
In the design of physical plausibility metrics, we refer to the design of the Skating Ratio in Gmd [29], as well as the designs for Floating, Ground Penetration, and Foot Clipping in ReinDiffuse [18], as well as the Speed and Smoothness metrics provided by Fischer [10]. The specific definitions of these metrics are as follows:
  • Skating Ratio: Foot skating refers to the situation where a person’s foot is in contact with the ground but still has velocity. Given a human motion $\Theta$, the Foot Skating Ratio for the $i$-th frame can be expressed as:
    $FS^{i}(\Theta)=\exp\left(-||\dot{\Theta^{i}_{ft}}\cdot f^{i}_{S}\cdot f^{i}_{V}\cdot f^{i}_{\bar{V}}||_{2}\right)$ (14)
    where $\dot{\Theta^{i}_{ft}}$, $f^{i}_{S}$, $f^{i}_{V}$, and $f^{i}_{\bar{V}}$ represent the foot velocity, the contact label, the instantaneous velocity threshold of the foot, and the average velocity threshold of the foot at the $i$-th frame, respectively.
  • Floating: Floating refers to a scenario where the lowest joint of the generated motion is not in contact with the ground. Specifically, the Floating for the $i$-th frame can be represented as:
    $Float^{i}(\Theta)=\exp\left(-||(\Theta^{i}_{h}-h_{ground})\cdot f^{i}_{F}||_{2}\right)$ (15)
    where $\Theta^{i}_{h}$ denotes the y-axis coordinate of the lowest point of the action at the $i$-th frame, $h_{ground}$ is a threshold hyperparameter, set to 5cm in practice, and $f^{i}_{F}$ is a flag that equals 1 when $\Theta^{i}_{h}>h_{ground}$ and 0 otherwise.
  • Ground Penetration: Ground Penetration refers to the situation where the lowest joint of the generated motion penetrates below the ground. The Ground Penetration for the $i$-th frame can be expressed as:
    $GP^{i}(\Theta)=\exp\left(-||(h_{ground}-\Theta^{i}_{h})\cdot f^{i}_{P}||_{2}\right)$ (16)
    where $f^{i}_{P}$ is a flag that equals 1 when $\Theta^{i}_{h}<h_{ground}$ and 0 otherwise.
  • Foot Clipping: Foot Clipping refers to a situation where the two feet are too close together, causing mesh penetration. The Foot Clipping for the $i$-th frame can be expressed as:
    $FC^{i}(\Theta)=\exp\left(-||(\Theta^{i}_{lf}-\Theta^{i}_{rf})\cdot f^{i}_{C}||_{2}\right)$ (17)
    where $\Theta^{i}_{lf}$ and $\Theta^{i}_{rf}$ are the joint coordinates of the left and right feet, and $f^{i}_{C}$ is a threshold function. $f^{i}_{C}=1$ when the distance between the two feet exceeds a threshold, which is set to 5cm in practice.
  • Speed: Speed can be used to calculate the amplitude of human motion. The calculation for Speed at the $i$-th frame is given by:
    $Speed^{i}(\Theta)=\frac{||\Theta^{i}||_{2}}{\Delta_{t}}$ (18)
    where $\Delta_{t}$ is the sampling frame rate of the dataset, which is 0.05 for HumanML3D and 0.08 for KIT-ML.
  • Smoothness: Smoothness can be used to measure the jittering of the model’s acceleration. Models with poor quality tend to exhibit poor Smoothness. The calculation for Smoothness at the $i$-th frame is given by:
    $Smooth^{i}(\Theta)=\frac{|\ddot{\Theta^{i}}-\ddot{\Theta^{i-1}}|}{\Delta_{t}}$ (19)
It is important to note that all of the aforementioned metrics should ideally be as close as possible to the dataset. Since the dataset includes motions such as crawling, climbing stairs, and freestyle swimming, metrics like Skating Ratio and Floating are not necessarily better when smaller. For the KIT-ML dataset [51], Ground Penetration and Clipping are also not zero, which may be related to the quality of the KIT-ML dataset.
Appendix B Implementation Details
For training, we employ DDPM [20] with 1,000 denoising steps ($T=1,000$) and linearly varying variances $\beta_{t}$ from 0.0001 to 0.02 in the forward process. The model is trained using AdamW [40] with an initial learning rate of 0.0002 and a weight decay of 0.01 for 50,000 iterations during pre-training, followed by 10,000 iterations of fine-tuning with a batch size of 64 on an NVIDIA A40 GPU. The learning rate is reduced by a factor of 0.9 every 5,000 steps. Gradient norms are clipped to a maximum value of 1 during gradient descent. For class-conditional generation (CFG) [19], text conditioning $c$ is set to $\emptyset$ for 10% of the samples, and the physical network is ignored with a probability of $P_{c}=10\%$. Additionally, we apply Exponential Moving Average (EMA) to smooth the parameter updates and reduce mutations and oscillations.
For inference, we utilize the SDE variant of the second-order DPM-Solver++ [41] with Karras sigmas [28] for 10-step sampling. The scale for text conditioning CFG is set to 2.5.
In our implementation of pose refinement, we used WHAM [60] as the baseline. However, the official WHAM code repository only provides versions trained on five datasets: AMASS [44], InstaVariety [27], MPI-INF-3DHP [45], Human3.6M [23], and 3DPW [70]. It does not include a version that incorporates the BEDLAM [4] dataset for training. As a result, the baseline results in our work differ from those presented in the WHAM paper.
Appendix C Network Architecture
Table 7: Model Architecture of BioMoDiffuse on dataset HumanML3D.
Model Module Details
BioMo Diffuse Physical Network (1-1)
• GRU layer
• AccelNet:
   - Motion MLP
   - Muscle MLP
   - Physics Decoder
   - Muscle Decoder
   - M_Net
   - F_Net
   - C_Net
   - Pose2EMG
Cond Unet1D (1-2)
• Time MLP
• Downsampling Blocks (×4) :
   - Conv1D
   - Cross Attention
   - Residual Blocks
• Mid Blocks (×2) :
• Upsampling Blocks (×4) :
   - Transposed Conv1D
   - Residual Blocks
Tab. 7 illustrates the network architecture of BioMoDiffuse trained on HumanML3D [14]. Compared to the baseline UNet, the addition of our physical network only increases the parameter count by $6.68\%$. However, this results in a significant improvement in the FID score, which decreases by $48.8\%$ (from 0.129 to 0.068). Other metrics also show notable improvements, demonstrating that the proposed physical network is both lightweight and effective.
Appendix D Derivation of Euler-Lagrange equation
The classical Euler-Lagrange equation is formulated as:
$\frac{d}{dt}\left(\frac{\partial L}{\partial\dot{q}}\right)-\frac{\partial L}{\partial q}=Q,$ (20)
where $L=T-V$ is the Lagrangian, $T$ denotes kinetic energy, $V$ represents potential energy, $q$ are generalized coordinates, and $Q$ encompasses non-conservative forces such as joint torques and external contact forces.
By modeling the human body as a rigid-body system, its kinetic energy can be expressed as:
$T=\frac{1}{2}\dot{q}^{T}M(q)\dot{q},$ (21)
where $M(q)$ is the inertia matrix. The potential energy $V(q)$ is defined by the gravitational field. Substituting $L=T-V$ into the Euler-Lagrange equation yields:
$\frac{d}{dt}\left(\frac{\partial T}{\partial\dot{q}}\right)-\frac{\partial T}{\partial q}+\frac{\partial V}{\partial q}=Q.$ (22)
Substituting the explicit form of $T$, we derive:
$M(q)\ddot{q}+\dot{M}(q)\dot{q}-\frac{1}{2}\dot{q}^{T}\frac{\partial M}{\partial q}\dot{q}+G(q)=Q,$ (23)
where $G(q)=\frac{\partial V}{\partial q}$. Leveraging the symmetry of $M(q)$, we reorganize indices and employ Christoffel symbols to define the Coriolis matrix:
$C(q,\dot{q})\dot{q}=\sum_{i,j}\frac{1}{2}\left(\frac{\partial M_{kj}}{\partial q_{i}}+\frac{\partial M_{ki}}{\partial q_{j}}-\frac{\partial M_{ij}}{\partial q_{k}}\right)\dot{q}_{i}\dot{q}_{j}.$ (24)
It can be rigorously shown that $C(q,\dot{q})\dot{q}=\dot{M}(q)\dot{q}-\frac{1}{2}\dot{q}^{T}\frac{\partial M}{\partial q}$, demonstrating that the Coriolis term arises from the derivative of the inertia matrix and quadratic velocity coupling, which depend solely on the body’s pose.
Figure 6: More samples from our BioMoDiffuse for text-to-motion synthesis. Samples generated with text prompts of HumanML3D’s test set. We recommend the supplemental video to see more motion results.
Appendix E Comparison on Samplers
Table 8: Comparative Evaluation of Motion Generation Methods. We compare different methods on generation quality, efficiency, and physical plausibility metrics using the HumanML3D dataset.
[…]
GT - - $0.002$ $2.974$ $0.797$ $0.124$ $0.682$ $0.358$ $2.60$ $0.205$ $0.057$ $0.000$ $0.000$
DEIS $10$ $3.659$ $0.148$ $2.732$ $0.843$ $0.298$ $2.388$ $0.380$ $0.91$ $0.195$ $0.054$ $0.013$ $0.002$
DEIS w/o Karras $10$ $3.615$ $0.146$ $2.723$ $0.844$ $0.271$ $2.285$ $0.381$ $0.88$ $0.195$ $0.051$ $0.013$ $0.002$
DPMSolver++ w/o Karras $10$ $3.664$ $0.126$ $2.755$ $0.840$ $0.147$ $1.071$ $0.374$ $1.39$ $0.196$ $0.050$ $0.012$ $0.001$
DPMSolver++ (20) $20$ $7.610$ $0.090$ $2.781$ $0.834$ $0.143$ $0.969$ $0.382$ $1.44$ $0.203$ $0.059$ $0.013$ $0.003$
DPMSolver++ (10) $10$ $2.553$ $0.071$ $2.784$ $0.835$ $0.135$ $0.959$ $0.385$ $1.45$ $0.204$ $0.056$ $0.013$ $0.002$
DDIM $50$ $19.82$ $0.123$ $2.741$ $0.839$ $0.132$ $0.933$ $0.368$ $1.57$ $0.185$ $0.054$ $0.012$ $0.002$
DDPM $1000$ $290.30$ $0.088$ $2.771$ $0.833$ $0.147$ $0.981$ $0.385$ $1.43$ $0.209$ $0.060$ $0.013$ $0.003$
In this section, we present the impact of different samplers on the generation performance of BioMoDiffuse. We evaluated the following samplers on our model: DDPM [20] , DDIM [62], DPMSolver [41], and DEIS [84]. Additionally, we incorporated Karras Sigma into DEIS and DPMSolver. Karras Sigma corresponds to constant-velocity thermal diffusion, which has been shown to ensure faster and higher-quality generation in the field of image synthesis. As shown in Tab. 8, we observe that DEIS fails to generate sufficiently smooth motions, while among the other samplers, DPMSolver achieves superior results in terms of FID and physical plausibility metrics. Furthermore, the inclusion of Karras Sigma contributes to improved FID scores. Based on these findings, we ultimately selected DPMSolver with Karras Sigma as the preferred configuration for our framework.
Appendix F User Study
Figure 7: User study interface for evaluating StableMoFusion and our model on HumanML3D.
In the user study, we generated multiple samples using the evaluation code on the HumanML3D test set, filtering out semantically redundant examples to create a final set of 20 distinct samples. Participants were asked to rate the generated motions on a 1–5 scale across four dimensions: ”Generation Quality,” ”Motion Smoothness,” ”Physical Plausibility,” and ”Text Alignment.” We invited 30 users to compare the motion generation results of our model, StableMoFusion, and Ground Truth based on these 20 prompts. To ensure unbiased feedback, the samples were presented in random order, preventing participants from identifying which motions were generated by our model.
Appendix G More Qualitative Results
Fig. 6 demonstrates the generation performance of our model across a broader set of prompts, while Fig. 8 provides a comparative analysis of our model against state-of-the-art methods on an extended set of prompts. The results indicate that our model consistently achieves superior performance across a wider range of motion types.
Figure 8: The comparison of the state-of-the-art methods on text-guided motion synthesis task. Samples generated with text prompts of HumanML3D’s test set. We recommend the supplemental video to see more motion results.

### v2

@@
-BioMoDiffuse: Physics-Guided Biomechanical Diffusion for Controllable and Authentic Human Motion Synthesis
+Biomechanics-Guided Residual Approach to Generalizable Human Motion Generation and Estimation
 Abstract
-Human motion generation holds significant promise in fields such as animation, film production, and robotics.
-However, existing methods often fail to produce physically plausible movements that adhere to biomechanical principles.
-While recent autoregressive and diffusion models have improved visual quality, they frequently overlook essential biodynamic features, such as muscle activation patterns and joint coordination, leading to motions that either violate physical laws or lack controllability.
-This paper introduces BioMoDiffuse, a novel biomechanics-aware diffusion framework that addresses these limitations.
-It features three key innovations: (1) A lightweight biodynamic network that integrates muscle electromyography (EMG) signals and kinematic features with acceleration constraints, (2) A physics-guided diffusion process that incorporates real-time biomechanical verification via modified Euler-Lagrange equations, and (3) A decoupled control mechanism that allows independent regulation of motion speed and semantic context.
-We also propose a set of comprehensive evaluation protocols that combines traditional metrics (FID, R-precision, etc.) with new biomechanical criteria (smoothness, foot sliding, floating, etc.).
-Our approach bridges the gap between data-driven motion synthesis and biomechanical authenticity, establishing new benchmarks for physically accurate motion generation.
-Figure 1: Our model, BioMoDiffuse, generates physically realistic motions based on textual input.
-By embedding Euler-Lagrange dynamics into the generation process, the model produces physical-optimized trajectories that inherently satisfy physical laws, achieving superior motion authenticity compared to current SOTA approaches.
-Specifically, BiMoDiffuse mitigates common kinematic artifacts including foot sliding, body interpenetration, and motion discontinuities while ensuring temporal smoothness.
-1 Introduction
-Human motion generation has wide-ranging applications in animation, film, and robotics.
-In recent years, there has been growing focus on generating human motion conditioned on language, sound, and motion trajectories [68, 61, 12, 49].
-Among these, language-guided motion generation stands out due to its high practical value, complexity, and the extensive foundational research supporting the field.
-Recent advancements [15, 48, 66, 17, 22, 24, 50, 88] in autoregressive and diffusion models have significantly improved the synthesis of diverse motion sequences while maintaining semantic consistency.
-While existing methods have made strides in enhancing the visual plausibility of motion sequences, studies in biomechanics and animation highlight significant shortcomings[54, 10] of them in capturing key aspects of human motion.
-The absence of explicit modeling for biomechanical and dynamic features prevents current models from learning essential information, such as acceleration patterns and muscle activation levels.
-Consequently, generated motions are often rigid and physically unrealistic, deviating significantly from natural human movement in rhythm and joint coordination.
-Additionally, current evaluation metrics fail to quantitatively assess physical plausibility and motion velocity, while existing methods struggle with precise control over movement amplitude and speed.
-These limitations considerably diminish the practical value of motion generation in professional applications that require biomechanical accuracy.
-To address these challenges, we propose BioMoDiffuse, a biomechanical-aware motion generation framework that seamlessly integrates musculoskeletal dynamics with diffusion processes.
-Our key innovation lies in embedding explicit physical supervision into the diffusion pipeline, leveraging a physics-aware network based on modified Euler-Lagrange equations and physics guidance to enforce biomechanical plausibility.
-Specifically, BioMoDiffuse introduces real-time biomechanical verification during denoising and a decoupled motion control system for independent regulation of velocity and semantics.
-Additionally, since current evaluation protocols neglect biomechanical-related aspects, we establish a set of comprehensive evaluation metrics emphasizing physical realism, combining biomechanical validity metrics (e.g., motion smoothness, foot sliding rate) with traditional generation metrics (FID, Diversity).
-Experiments on HumanML3D [14] and KIT-ML [51] demonstrate significant improvements over existing methods.
-The main contributions of this paper are as follows:
-  • A biomechanics-guided diffusion framework that integrates EMG signals and kinematic features with acceleration-based physical supervision, ensuring compliance with rigid-body dynamics .
-  • A physics-aware network architecture based on Euler-Lagrange principles, incorporating a novel motion control system for enhancing controllability.
-  • A comprehensive evaluation framework that emphasizes physical realism, unifying biomechanical metrics and traditional metrics.
-2 Related Work
-Physical plausible motion processing.
+Human pose, action, and motion generation are critical for applications in digital humans, character animation, and humanoid robotics.
+However, many existing methods struggle to produce physically plausible movements that are consistent with biomechanical principles.
+Although recent autoregressive and diffusion models deliver impressive visual quality, they often neglect key biodynamic features and fail to ensure physically realistic motions.
+Reinforcement Learning (RL) approaches can address these shortcomings but are highly dependent on simulation environments, limiting their generalizability.
+To overcome these challenges, we propose BioVAE, a biomechanics-aware framework with three core innovations: (1) integration of muscle electromyography (EMG) signals and kinematic features with acceleration constraints to enable physically plausible motion without simulations; (2) seamless coupling with diffusion models for stable end-to-end training; and (3) biomechanical priors that promote strong generalization across diverse motion generation and estimation tasks.
+Extensive experiments demonstrate that BioVAE achieves state-of-the-art performance on multiple benchmarks, bridging the gap between data-driven motion synthesis and biomechanical authenticity while setting new standards for physically accurate motion generation and pose estimation.
+Index Terms:
+Human Motion Generation, Pose Estimation, Diffusion Models.
+(a) Results across multiple tasks
+(b) Quantitative comparison
+Figure 1: (a) Demonstrates our method’s results across multiple tasks, including motion generation, motion refinement, pose estimation, and motion editing.
+(b) Shows a quantitative comparison of different metrics against current state-of-the-art methods.
+Our proposed BioVAE framework demonstrates strong generalization across these tasks.
+Compared to state-of-the-art methods, our approach generates motions with significantly reduced foot floating, penetration, and clipping artifacts, as well as improved temporal smoothness.
+Notably, BioVAE achieves a clear improvement in physical plausibility metrics, setting a new benchmark for biomechanically realistic motion synthesis.
+footnotetext: Zixi Kang, Xinghan Wang and Yadong Mu are with the Wangxuan Institute of Computer Technology, Peking University, Beijing 100871, China (e-mail: forever.kzx0713@stu.pku.edu.cn, xinghan_wang@pku.edu.cn, myd@pku.edu.cn).
+Yadong Mu is the corresponding author.
+I Introduction
+Human pose, action, and motion generation have wide-ranging applications in motion capture, animation, and robotics.
+In recent years, there has been a growing focus on generating human motion conditioned on image, language, sound, and motion trajectories [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
+However, studies in biomechanics and animation[13, 14] have shown that current methods fail to capture the distinct characteristics of human motion, resulting in generated sequences that often deviate significantly from natural movement, with unnatural rhythms and unrealistic joint configurations.
+Moreover, the lack of explicit modeling of biodynamic features prevents these models from capturing critical information such as acceleration patterns and muscle activation levels, often resulting in motions that appear stiff or violate physical laws.
+Furthermore, existing evaluation metrics also overlook the quantitative assessment of physical plausibility and motion velocity, while current approaches struggle to control movement amplitude and speed precisely.
+These limitations severely undermine the practical value of motion generation in professional applications that demand high biomechanical precision.
+To address these challenges, we propose BioVAE, a biomechanics-aware framework that seamlessly integrates musculoskeletal dynamics into diffusion processes.
+Our key innovation lies in embedding Euler-Lagrange equations and EMG signals into the VAE architecture, while connecting the VAE residuals to the diffusion process to enable end-to-end, generalizable motion generation.
+Furthermore, since current evaluation protocols largely overlook biomechanical aspects, we establish a comprehensive set of metrics which emphasizes physical realism, combining biomechanical validity measures (e.g., motion smoothness, foot sliding rate) with traditional generation metrics such as FID and diversity.
+Extensive experiments on HumanML3D[15], KIT-ML[16], 3DPW[17], and RICH[18] demonstrate that BioVAE achieves significant improvements over existing methods.
+The main contributions of this paper are as follows.
+  • We propose BioVAE, a residual approach based on Euler-Lagrange equations that can be jointly trained with diffusion models in a self-supervised setting.
+  • We introduce EMG signals and acceleration-based physical supervision, leveraging additional biomechanical priors to enhance the model’s capability.
+  • We design extensive experiments across different tasks to evaluate both reconstruction quality and physical plausibility, demonstrating the strong generalizability of our framework.
+II Related Work
+II-A Physically plausible motion processing.
 In motion generation, Reinforcement Learning (RL) has been utilized to fine-tune models for physical plausibility [19, 20, 21].
 For example, ReinDiffuse [22] and PhysDiff [23] extend MDM [24] to improve motion quality.
-However, these RL-based approaches predominantly focus on refining foot-ground contact using opaque, black-box reward functions, often overlooking broader rigid-body dynamics and biomechanical principles.
-The lack of transparency in their loss design further constrains interpretability and limits precise control over kinematic behaviors.
-Alternatively, some methods incorporate auxiliary constraints during training [75, 29, 53, 34]; for instance, OmniControl [75] leverages spatial signals to guide generation.
-Nonetheless, these approaches tend to prioritize controllability at the expense of enforcing comprehensive physical laws, disproportionately emphasizing foot-centric adjustments rather than ensuring holistic motion plausibility.
-Physical plausibility is also addressed in other motion-related tasks that demand biomechanically authentic synthesized motion sequences, such as human pose estimation from videos [32, 37, 33, 63, 60].
-Recent advances try to incorporate physics-based modeling [78, 58, 87, 86, 59, 52, 35, 74] to enhance the quality of the estimated motion sequences.
-However, these methods still face challenges of suboptimal precision and excessive jitter.
-Although techniques like Gaussian filtering can suppress high-frequency noise, they often compromise precision.
-Consequently, several studies [76, 7, 25, 79, 69, 30, 65] have explored pose refinement techniques to reduce temporal instability.
-For instance, SmoothNet [79] employs a parametric approach to optimize the temporal dimension, substantially enhancing motion quality.
-Diffusion models.
-Diffusion models have demonstrated remarkable success in image generation, with various methods leveraging language guidance to produce high-quality images [55, 57, 81].
-Among these approaches, DDPM [20] has emerged as the most widely adopted sampler, achieving impressive results.
-Building on DDPM, DDIM [62] reduces the number of inference steps while enabling deterministic sampling by adjusting the variance.
-The introduction of Classifier Guidance (CG) [9] empowered diffusion models to handle conditional generation tasks effectively, and CFG [19] further streamlined this process by replacing CG, enhancing the efficiency of guidance integration during training.
-In this paper, we observe that incorporating additional domain knowledge parallels the mechanisms of CG and CFG conceptually, leading us to re-derive the diffusion process accordingly.
-In recent years, many studies have tackled the Diffusion Inversion problem, which provides a method for mapping real images to latent space, with broad applications in image editing [46, 71, 36, 26, 85, 8].
-However, Diffusion Inversion methods have rarely been applied to tasks outside of image generation.
-We observe that both motion refinement and image editing aim to preserve the macroscopic structure while modifying finer details.
-Consequently, we are the first to apply DDIM Inversion to the task of motion refinement.
-Text-driven human motion generation.
+However, these RL-based methods suffer from two major limitations: (1) they lack the ability of generalization and require task-specific, dataset-specific, and model-specific reward designs; and (2) they depend heavily on simulation environments, making the training process a black box and preventing end-to-end optimization.
+Alternatively, some methods incorporate auxiliary constraints during training [25, 26, 27]; for instance, OmniControl [25] leverages spatial signals to guide generation.
+However, since these approaches mainly focus on enabling multimodal and editable motion generation instead of improving motion quality, they remain prone to challenges like limited precision and excessive jitter.
+Although techniques like Gaussian filtering can suppress high-frequency noise, they often degrade motion precision.
+Consequently, several studies [28, 29, 30, 31, 32, 33, 34, 35] have explored pose refinement techniques to reduce temporal instability.
+For example, SmoothNet [31] introduces a parametric approach to optimize temporal coherence and substantially improve motion quality.
+Nevertheless, these methods do not address generalization and cannot be seamlessly integrated with diffusion models or large language models (LLMs), failing to leverage the powerful generative capabilities of such backbones and exhibiting inherent limitations.
+II-B Diffusion models.
+Diffusion models have achieved remarkable success in image generation, with numerous approaches leveraging language guidance to produce high-quality results [36, 37, 38].
+Among them, Denoising Diffusion Probablistic Models (DDPM) [39] has become the most widely adopted sampling method, delivering impressive results.
+Based on DDPM, DDIM [40] reduces inference steps and allows deterministic sampling by adjusting the variance schedule.
+The introduction of Classifier Guidance (CG)[41] extended diffusion models to conditional generation tasks, while Classifier-Free Guidance (CFG)[42] further streamlined this process by replacing CG and improving the efficiency of the guidance during training.
+Inspired by these advancements, we observe that incorporating domain-specific knowledge is conceptually aligned with the mechanisms of CG and CFG, thereby motivating a re-derivation of the diffusion process.
+In parallel, recent studies have addressed the diffusion inversion problem, which map real images to latent space and supports various image editing applications [43, 44, 45, 46, 47, 48].
+However, diffusion inversion techniques have rarely been explored beyond image generation.
+This work is the first to apply DDIM inversion to motion generation tasks.
+II-C Text-driven human motion generation.
 Early approaches to text-driven human motion generation conceptualized motion as a form of language, adopting autoregressive methods commonly used in language models [49, 50, 51, 52].
-Recognizing the potential of large language models, subsequent studies improved motion embeddings to enable the joint learning of motion and text, yielding promising results [15, 16, 80, 11, 73, 38, 42].
+Recognizing the potential of large language models, subsequent studies improved motion embeddings to enable joint learning of motion and text, yielding promising results [53, 54, 55, 56, 57, 58].
 MDM [24] pioneered the application of diffusion models to human motion generation, while MLD [59], inspired by VAE [60], introduced a diffusion process within the latent space.
@@
 MoFusion [63] and StableMoFusion [64], drawing inspiration from Stable Diffusion [36], utilized the UNet [65] architecture without relying on external knowledge, achieving state-of-the-art results.
-Notably, experiments with StableMoFusion revealed that DiT [47] and RetNet [64] underperformed relative to UNet, validating our choice of UNet as the backbone for the diffusion framework.
-3 Method
-Current approaches fail to capture the essential characteristics of human motion due to the absence of biomechanical constraints and rigid-body dynamics modeling, limiting their ability to learn authentic movement patterns.
-To address this gap, we present a novel motion generation framework that explicitly incorporates biomechanical principles and rigid-body kinematics, which is capable of synthesizing authentic physically plausible human motions.
-Our methodological contributions are as follows: Sec.
-3.1 establishes the theoretical foundation by formalizing the biomechanical constraints; Sec.
-3.2 introduces a physics-aware network architecture derived from Euler-Lagrange equations, ensuring dynamic compliance; Sec.
-3.3 presents our biomechanics-guided diffusion based on joint probability distribution analysis.
-The complete motion generation framework is detailed in Sec.
-3.4.
-Furthermore, Sec.
-3.5 demonstrates that our proposed method, equipped with diffusion inversion, can be effectively utilized for physics-aware motion refinement.
-Figure 2: Overview of BioMoDiffuse.
-At each diffusion timestep $t$, our framework processes the motion sequence $\Theta^{t}$ alongside text conditions.
-The text is encoded using a frozen CLIP encoder and concatenated with $\Theta^{t}$ as input to the UNet backbone.
-The light brown panel (left) illustrates our physics-aware network, which consists of: (1) An Accel Net that implements augmented Euler-Lagrange equations to predict biomechanically valid accelerations, and (2) An Autoregressive Block that models temporal dependencies.
-The training objectives combine: $\mathcal{L}_{\text{phys}}$ (from Accel Net), $\mathcal{L}_{\text{diffusion}}$ (reconstruction), and $\mathcal{L}_{\text{accel}}$ (acceleration supervision against ground truth).
-Structural details of the orange (autoregressive block) and light blue (Accel Net) components are provided in the right panels.
-3.1 Preliminaries
-Text-guided human motion generation requires an input text to condition the generation of a motion sequence $\Theta\in\mathbb{R}^{T\times D}$, where $T$ denotes the sequence length and $D$ represents the dimension of the joint representation.
-This joint representation can be converted into a coordinate representation $\Phi\in\mathbb{R}^{T\times 3d}$, where $d$ is the number of joints.
-The Euler-Lagrange equation is given by:
-$F=M\cdot\ddot{\Phi_{t}}+C,$ (1)
+To enable our BioVAE to deeply integrate with the diffusion process, we formulate BioVAE as a self-supervised residual connection, re-derive the probability formulation of the diffusion process, and redesign the loss function to achieve stable end-to-end training of the two components.
+III Method
+Figure 2: Overview of BioVAE.
+We employ the MIA dataset [66] and adopt a training strategy that learns the mapping from motion sequences to EMG signals within the motion encoder.
+In the latent space, a neural network predicts four intermediate variables: $M^{-1}$, $E$, $C$, and $A$, which are then combined through the Euler-Lagrange equations to reconstruct the joint local rotational accelerations.
+The motion decoder iteratively generates residuals from the acceleration information to guide the UNet’s generation process.
+BioVAE can be applied to a variety of tasks and modalities, such as text-to-motion, pose estimation, motion editing, and motion refinement, demonstrating strong generalization across tasks while significantly enhancing the physical plausibility of the generated motions.
+Existing methods fail to capture the unique characteristics of human motion due to the absence of biomechanical constraints and the lack of rigid body modeling.
+In this section, we present our BioVAE framework, which integrates biomechanical principles and rigid body dynamics into the network architecture, providing a diffusion-based model that achieves state-of-the-art performance in multiple tasks.
+We begin in Section III-A by introducing the preliminaries.
+Section III-B details the architecture of the BioVAE network.
+In Section III-C, we derive the joint distribution formulation that allows end-to-end training of BioVAE with diffusion through residual connections.
+Finally, Section III-D describes our diffusion inversion strategy, showing how BioVAE supports motion refinement, motion editing, and pose estimation.
+III-A Preliminaries
+Human motion generation produces a motion sequence $\Theta\in\mathbb{R}^{T\times D}$, where $T$ denotes the length of the sequence and $D$ is the dimension of the joint representation.
+This motion sequence can be transformed into a joint coordinate representation $X\in\mathbb{R}^{T\times 3d}$, where $d$ represents the number of joint nodes.
+By modeling the human body as an articulated rigid-body system, the Euler-Lagrange equation can be formulated as:
+$F=M\cdot\ddot{X_{t}}+C,$ (1)
 where $F$ represents the internal and external contact forces, $M$ is the inertia matrix, and $C$ captures the combined effect of the Coriolis force and gravity.
@@
 where $R$ is the muscle rotation matrix, $A(u)$ is the muscle activation force, and $u$ is the muscle activation coefficient.
-Using this decomposition, the Euler-Lagrange equation for human motion can be rewritten as Eq.
-3.
-Please refer to Appendix D for details.
-$\ddot{\Phi_{t}}=M^{-1}(E+C+R\cdot A(u)).$ (3)
-3.2 Physics-Aware Network
-The Euler-Lagrange equation-based network provides a stronger prior than conventional neural networks, as it inherently captures acceleration dynamics while reducing problem complexity.
-A crucial design here is to directly embed the constraints within the network architecture itself.
-In a physical sense, the Euler-Lagrange equation serves as a rigorous constraint for human motion: by framing motion generation as $Motion=Gen(text)$, this equation ensures the learned function aligns closely with real-world biomechanics; From a deep learning perspective, Euler-Lagrange equation networks offer a more robust alternative to traditional non-parametric constraints for handling diverse human skeleton data in supervised settings.
-Additionally, the inherent properties of Euler-Lagrange equations in our problem context guarantee convergence to a stable and unique solution.
-Following the MIA dataset representation, a learnable mapping exists between human poses and the activation coefficients of primary muscles $u\in\mathbb{R}^{8}$, defined as $u=u(\Phi_{:t})$.
-The muscle rotation matrix $R$ depends solely on pose rotation, while the inertia matrix $M$—a symmetric matrix—is intrinsic to the body and determined by position.
-The Coriolis force and gravity are formulated as:
-$C=\dot{M}(\Phi_{t})\dot{\Phi_{t}}-\frac{1}{2}\dot{\Phi_{t}}^{T}\frac{\partial M}{\partial\Phi_{t}}\dot{\Phi_{t}}+G(\Phi_{t}),$ (4)
-which are also pose-dependent. External contact forces $F$ (e.g., elastic/resistive forces) depend on both position and velocity. Thus, Eq. 3 is reformulated as:
-$\ddot{\Phi}_{t}=M^{-1}(\Phi_{:t})\left(F(\Phi_{:t})+C(\Phi_{:t})+R(\Phi_{:t})\odot A(u)\right),$ (5)
-where $\odot$ denotes element-wise multiplication.
-We implement three neural networks to approximate $M$, $F$, and $C$, with $M$ constrained as symmetric.
-The term $R\odot A(u)$ is modeled via cross-attention, enabling lightweight networks to predict pose acceleration.
-Finally, leveraging autoregressive processing, the predicted $\ddot{\Phi}_{t}$ and speed control signal are integrated as inputs to derive the optimized $\Theta_{t}$, ensuring temporal coherence and biomechanical fidelity.
-3.3 Physics Guidance
-We introduce additional guidance into the diffusion process to ensure that acceleration remains physically meaningful and achieve more stable training and faster convergence.
-Our network explicitly treats the acceleration $\ddot{\Phi_{t}}$ as an intermediate variable.
-We aim for the acceleration $\ddot{\Phi_{t}}$ in our physics-based network to converge to the real result as quickly as possible.
-To achieve this, we impose a supervised constraint on the acceleration during the diffusion process.
-Specifically, at each reverse diffusion step, we consider extending the distribution at step $t$, denoted by $\Theta^{t}$, into a joint distribution $S^{t}=\{\Theta^{t},\ddot{\Phi}^{t}\}$.
-According to DDPM theory, when performing diffusion on $S^{t}$, the model’s loss function can be written as:
-$\min E_{Q(S^{0}:S^{T})}\left[-\log P_{\theta}(S^{0}|S^{1})\right.$ (6)
-$+\sum_{t=2}^{T}\text{KL}(Q(S^{t-1}|S^{t},S^{0})\|\left.P_{\theta}(S^{t-1}|S^{t}))\right].$
-We argue that, during the denoising process, the quantities $\Theta^{t-1},\ddot{\Phi}^{t-1}$ at step $t-1$ are unknown and do not satisfy the interpolation relationship.
-However, the known quantities $\Theta^{t},\ddot{\Phi}^{t}$ at step $t$ do satisfy this relationship.
-This leads to the following:
-$\log P_{\theta}(S^{t-1}|S^{t})=\log P_{\theta}(\Theta^{t-1},\ddot{\Phi}^{t-1}|\Theta^{t},\ddot{\Phi}^{t})$ (7)
-$=\log P_{\theta}(\Theta^{t-1},\ddot{\Phi}^{t-1}|\Theta^{t})$
-$=\log P_{\theta}(\Theta^{t-1}|\ddot{\Phi}^{t-1},\Theta^{t})+\log P_{\theta}(\ddot{\Phi}^{t-1}|\Theta^{t}).$
-In the expression in Eq.
-7, $\log P_{\theta}(\Theta^{t-1}|\ddot{\Phi}^{t-1},\Theta^{t})$ resembles the condition $c$ in CG and CFG, while $\log P_{\theta}(\ddot{\Phi}^{t-1}|\Theta^{t})$ can be obtained using the physical network we propose.
-Thus, during training, we omit the physical network with probability $P_{c}$, and use it with probability $1-P_{c}$.
-We set $\ddot{\Phi}^{t-1}=\text{PhysNet}(\Theta^{t})$, and the predicted acceleration is compared to the ground truth acceleration, introducing a physics loss $L_{\text{phys}}=\text{MSE}(\text{PhysNet}(\Theta^{t}),\ddot{\Phi_{\text{gt}}}^{t})$.
-To further enhance acceleration accuracy, we design the following additional loss term for supervision:
-$L_{\text{accel}}=\text{MSE}(\ddot{\Phi}^{t},\ddot{\Phi_{\text{gt}}}^{t}).$ (8)
-Thus, the total loss function during training can be written as:
-$L=L_{\text{diffusion}}+L_{\text{accel}}+L_{\text{phys}}.$ (9)
-3.4 Model Architecture
-Table 1: Quantitative evaluation on the HumanML3D and KIT-ML test set.
-$\pm$±indicates a 95% confidence interval.
+Using this decomposition, the Euler-Lagrange equation for human motion can be rewritten as Equation 3.
+$\ddot{X_{t}}=M^{-1}(E+C+R\cdot A(u)).$ (3)
+III-B BioVAE Architecture
+Compared to conventional neural networks, a network based on the Euler-Lagrange equation provides a stronger prior, as it can inherently capture acceleration dynamics while leveraging constraints to reduce the complexity of the problem.
+In this section, we introduce how to design a VAE architecture guided by the Euler-Lagrange equation.
+Following the representation in the MIA dataset, there exists a learnable mapping between human poses and the activation coefficients of the primary muscles $u\in\mathbb{R}^{8}$, defined as $u=u(X_{:t})$.
+The muscle rotation matrix $R$ depends solely on the rotational component of the pose, while the inertia matrix $M$ is an intrinsic property of the body determined only by position.
+The Coriolis force and gravity can be expressed as follows.
+$C=m\omega^{2}r+2m\omega\times\dot{X}_{t}-\nabla U_{gravity}(X_{t}),$ (4)
+which are likewise pose-dependent.
+The external contact force $F$, including the elastic and resistive forces, depends on both position and velocity.
+Using this formulation, Equation (3) can be rewritten as a function of $X_{:t}$ in Equation (5):
+$\ddot{X}_{t}=M^{-1}(X_{:t})\left(F(X_{:t})+C(X_{:t})+R(X_{:t})\cdot A(u)\right).$ (5)
+We implement three neural networks to approximate $M^{-1}$, $F$, and $C$, where $M^{-1}$ is restricted to be symmetric.
+The term $R\cdot A(u)$ is modeled using cross-attention, allowing four lightweight networks to predict the acceleration of the pose effectively.
+Once acceleration $\ddot{X}_{t}$ is predicted, we employ an autoregressive decoder to reconstruct the residual between predicted $X_{t}$ and the ground truth.
+This enables optimization at each training time step while ensuring temporal consistency and biomechanical fidelity.
+TABLE I: Quantitative evaluation on the HumanML3D and KIT-ML test set.
+The confidence interval is indicated by±$\pm$.
 Bold face indicates the best result, while underscore refers to the second best.
@@
 StableMoFusion $0.445^{\pm.006}$ $0.660^{\pm.005}$ $0.782^{\pm.004}$ $0.258^{\pm.029}$ $2.832^{\pm.014}$ $10.94^{\pm.077}$ $1.36^{\pm.062}$
+MoMask $0.433^{\pm.007}$ $0.656^{\pm.005}$ $0.781^{\pm.005}$ $0.204^{\pm.011}$ $2.779^{\pm.022}$ $1.13^{\pm.043}$
 Ours $0.448^{\pm.008}$ $0.666^{\pm.005}$ $0.788^{\pm.005}$ $0.211^{\pm.101}$ $2.772^{\pm.017}$ $11.11^{\pm.094}$ $1.38^{\pm.050}$
@@
 StableMoFusion $0.545^{\pm.003}$ $0.743^{\pm.003}$ $0.833^{\pm.002}$ $0.133^{\pm.005}$ $2.814^{\pm.008}$ $9.704^{\pm.065}$ $1.844^{\pm.060}$
+MoMask $0.521^{\pm.002}$ $0.713^{\pm.002}$ $0.807^{\pm.002}$ $\bm{0.045}^{\pm.002}$ $2.958^{\pm.008}$ $9.644^{\pm.086}$ $1.241^{\pm.040}$
 Ours $0.547^{\pm.003}$ $0.743^{\pm.002}$ $0.835^{\pm.002}$ $0.071^{\pm.003}$ $2.784^{\pm.008}$ $9.567^{\pm.086}$ $1.919^{\pm.063}$
-Our biomechanics-aware diffusion framework is built upon three core components, as shown in Fig. 2. The architecture seamlessly integrates biomechanical priors with diffusion dynamics through:
-Biomechanical Encoding.
-Using the MIA dataset [6], we pre-train a neuromuscular encoder that establishes a dynamical mapping $\mathcal{F}:\Theta\rightarrow u$ between joint angles $\Theta$ and muscle activation coefficients $u$.
-This encoder remains differentiable throughout end-to-end training.
-Physics-Conditioned Diffusion. The UNet-based diffusion backbone processes motion sequences $\mathbf{x}_{t}$ at timestep $t$ through:
-$z_{\theta}=\text{BioMoDiffuse}\big{(}\text{CLIP}(\mathbf{c}),\mathbf{x}_{t},\mathcal{F}(\Theta)\big{)},$ (10)
-where text conditioning $\mathbf{c}$ is incorporated via cross-attention layers.
-We adopt a two-phase training process: initial diffusion followed by physics-constrained fine-tuning (which introduces the physics loss $\mathcal{L}_{\text{phys}}$).
-Controllable Generation.
-Our decoupled control mechanism enables independent manipulation of motion dynamics and semantics through: (1) classifier-free guidance and physics guidance weights, and (2) the injection of biomechanical constraints via the pre-trained encoder.
-3.5 Pose Refinement with Diffusion Inversion
-Figure 3: Refining pose using our method.
-Building on existing pose estimation models, we integrate BioMoDiffuse with DDIM inversion [62].
-By adopting an inversion-then-forward pipeline, our approach significantly improves the smoothness of generated SMPL sequences while enhancing the fitting accuracy of pose estimation.
-Our proposed framework can be readily adapted to refine pose estimates from existing pose estimation methods.
-The key insight is that our physics-guided diffusion model is capable of generating physically plausible motions starting from latent noise.
-Given the initial estimates of human poses, we first employ diffusion inversion—a method that reverses the usual diffusion process—to recover the latent noise corresponding to the poses.
-We then feed this noise into the physics-guided diffusion model, which injects biomechanical constraints to produce the refined poses that are more authentic compared to original estimates.
-Specifically, given the initial SMPL [39] parameters $\mathbf{x}_{0}$ from pose estimation, we first invert them to timestep $t$ using DDIM sampling:
+III-C End-to-End Residual Training
+In previous approaches, optimization methods such as reinforcement learning (RL) or refinement [67, 31] are often applied as post-processing on top of the backbone model.
+In contrast, we integrate BioVAE with diffusion to enable end-to-end training.
+Furthermore, we apply optimization at every diffusion training timestep, allowing our model to perform iterative refinement during inference, which substantially improves the quality of generation.
+Traditional RL and simulation-based methods [21, 20, 23] typically require a complete human motion sequence and thus cannot be integrated into the noisy training process of diffusion models.
+To address this, our network explicitly treats acceleration $\ddot{X_{t}}$ as an intermediate variable.
+We introduce supervised constraints on the acceleration during the diffusion process to ensure that the predicted accelerations are physically meaningful, resulting in a more stable training process and faster convergence.
+Specifically, at each reverse diffusion step, we consider extending the distribution at step $t$, denoted as $\Theta^{t}$, to a joint distribution $S^{t}=\{\Theta^{t},\ddot{X}^{t}\}$.
+According to DDPM theory, when performing diffusion on $S^{t}$, the model’s loss function can be written as:
+$\min E_{Q(S^{0}:S^{T})}\left[-\log P_{\theta}(S^{0}|S^{1})\right.$ (6)
+$+\sum_{t=2}^{T}\text{KL}(Q(S^{t-1}|S^{t},S^{0})\|\left.P_{\theta}(S^{t-1}|S^{t}))\right].$
+We argue that, during the denoising process, the quantities $\Theta^{t-1},\ddot{X}^{t-1}$ at step $t-1$ are unknown and do not satisfy the interpolation relationship.
+However, the known quantities $\Theta^{t},\ddot{X}^{t}$ at step $t$ do satisfy this relationship.
+This leads to the following:
+$\log P_{\theta}(S^{t-1}|S^{t})=\log P_{\theta}(\Theta^{t-1},\ddot{X}^{t-1}|\Theta^{t},\ddot{X}^{t})$ (7)
+$=\log P_{\theta}(\Theta^{t-1},\ddot{X}^{t-1}|\Theta^{t})$
+$=\log P_{\theta}(\Theta^{t-1}|\ddot{X}^{t-1},\Theta^{t})+\log P_{\theta}(\ddot{X}^{t-1}|\Theta^{t}).$
+In the expression in Equation 7, $\log P_{\theta}(\Theta^{t-1}|\ddot{X}^{t-1},\Theta^{t})$ resembles the condition $c$ in CG and CFG, while $\log P_{\theta}(\ddot{X}^{t-1}|\Theta^{t})$ can be obtained using the VAE architecture we propose.
+Thus, during training, we set $\ddot{X}^{t-1}=\text{VAE}(\Theta^{t})$, and the predicted acceleration is compared to the ground truth acceleration, introducing a physics loss $L_{\text{phys}}=\text{MSE}(\text{VAE}(\Theta^{t}),\ddot{X_{\text{gt}}}^{t})$.
+To further enhance acceleration accuracy, we design the following additional loss term:
+$L_{\text{accel}}=\text{MSE}(\ddot{X}^{t},\ddot{X_{\text{gt}}}^{t}).$ (8)
+Thus, the total loss function during training is written as:
+$L=L_{\text{diffusion}}+L_{\text{accel}}+L_{\text{phys}}.$ (9)
+III-D Generation with Diffusion Inversion
+Figure 3: Estimation, refinement, and editing using DDIM Inversion.
+We incorporate the BioVAE-Diffusion framework as a residual module and apply diffusion inversion to enable pose estimation, motion refinement, and motion editing tasks.
+By adopting an inversion-then-forward pipeline, our approach significantly improves the smoothness of generated SMPL sequences while enhancing the overall generation quality.
+BioVAE can be readily adapted to pose estimation, motion refinement, and motion editing tasks.
+The key insight is that our BioVAE-Diffusion framework is capable of generating physically plausible motions starting from latent noise.
+Given the initial estimates of human motion, we first employ diffusion inversion [44, 46]—a technique that reverses the usual diffusion process—to recover the latent noise corresponding to the motion sequence.
+We then feed this noise into our model, which injects biomechanical constraints to produce motion sequences that are more authentic compared to the original inputs.
+Specifically, given the initial SMPL [68] parameters $\mathbf{x}_{0}$, we first invert them to timestep $t$ using DDIM sampling:
 $\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\epsilon_{\theta}(\mathbf{x}_{0},t),$ (10)
-where $t$ controls the balance between information preservation and noise introduction.
-The inverted latent $\mathbf{x}_{t}$ then undergoes forward diffusion with our physics guidance to produce the refined poses:
+where $t$ controls the balance between information preservation and noise introduction. The inverted latent $\mathbf{x}_{t}$ then undergoes forward BioVAE-Diffusion to produce the refined motion:
 $\hat{\mathbf{x}_{0}}=f_{\theta}(\mathbf{x}_{t},t).$ (11)
-Crucially, our parametric implementation inherits the physics constraints from Secs.
-3.2 and 3.3 without the need for additional tuning, demonstrating a unified biomechanical framework across both motion generation and refinement tasks.
-4 Experiments
-4.1 Datasets and Metrics
-We conducted training and testing on our model using both the KIT-ML [51] and HumanML3D [14] datasets.
-The KIT-ML dataset contains 3,911 motions and 6,363 natural language annotations, while HumanML3D is derived from and reannotates the HumanAct12 [13] and AMASS [44] datasets, comprising a total of 14,616 motion sequences and 44,970 descriptions.
-We follow the evaluation metrics used in previous motion generation works [66, 82, 22], including Frechet Inception Distance (FID), Precision, Diversity, Multimodality, and Multi-Modal Distance.
-Additionally, we consider physics-related metrics such as skating ratio, ground floating, ground penetration, and foot clipping, as used in prior studies [18, 75, 77].
-Furthermore, we introduce two novel metrics based on Fourier analysis: Velocity Consistency (VC) and Acceleration Consistency (AC).
-Our motivation stems from the observation that while the Smoothness metric captures global temporal coherence, it overlooks the inherent kinematic relationship between stable body parts and agile extremities in human motion.
-For a motion sequence represented as $\Phi\in\mathbb{R}^{T\times 3d}$, the metrics are formally defined as:
-VC $=\frac{1-\|\text{FFT}(\dot{\Phi})_{[:thre_{\text{low}}]}\|-\|\text{FFT}(\dot{\Phi})_{[thre_{\text{high}}:]}\|}{\|\text{FFT}(\dot{\Phi})_{[:thre_{\text{low}}]}\|+\|\text{FFT}(\dot{\Phi})_{[thre_{\text{high}}:]}\|},$ (13)
-AC $=\frac{1-\|\text{FFT}(\ddot{\Phi})_{[:thre_{\text{low}}]}\|-\|\text{FFT}(\ddot{\Phi})_{[thre_{\text{high}}:]}\|}{\|\text{FFT}(\ddot{\Phi})_{[:thre_{\text{low}}]}\|+\|\text{FFT}(\ddot{\Phi})_{[thre_{\text{high}}:]}\|}.$
-Spectral ratio analysis enables VC and AC to effectively characterize overall motion patterns by precisely quantifying the balance between high-frequency components (associated with rapid limb movements) and low-frequency components (correlated with stable torso dynamics).
-This frequency-domain approach provides a more accurate assessment of motion physicality compared to traditional smoothness metrics.
-Since the dataset includes motions like “climbing stairs” (a foot is floating from the ground) and “jumping while spinning” (abrupt acceleration), all these physics-based metrics should be made as consistent as possible with ground-truth instead of being blindly minimized to meet the characteristics of dataset.
-4.2 Quantitative Results
-Figure 4: User Study Results.
-User Study.
-We asked participants to rate four aspects of the generated motions on a 1–5 scale: “Generation Quality,” “Motion Smoothness,” “Physical Plausibility,” and “Text Alignment.” Using the HumanML3D test set.
-Thirty users were invited to compare the motion generation results of our model, StableMoFusion, and Ground Truth across 20 prompts.
-The results demonstrate that our model exhibits superior performance in terms of generation quality, motion smoothness, physical plausibility, and semantic alignment compared to StableMoFusion.
-Furthermore, on certain examples, our method achieves results that are comparable to the ground truth, showcasing its competitive performance.
-Table 2: Quantitative evaluation of physical plausible metrics on the HumanML3D and KIT-ML test set.
+Crucially, our parametric implementation inherits the physics constraints from Sections III-B and III-C without requiring additional tuning, demonstrating the strong generalization capability of our method across multiple tasks.
+TABLE II: Quantitative evaluation of physical plausible metrics on the HumanML3D and KIT-ML test set.
 All experiments were repeated 20 times with different random seeds, and the mean results are reported in the tables.
@@
 StableMoFusion $0.292$ $2.220$ $0.448$ $0.74$ $0.234$ $0.074$ $0.043$ $0.002$
-Ours $\bm{0.135}$ $\bm{0.959}$ $0.385$ $\bm{1.45}$ $\bm{0.204}$ $\bm{0.056}$ $\bm{0.013}$ $0.002$
+MoMask $0.106$ $1.048$ $0.331$ $\bm{1.89}$ $0.261$ $0.050$ $0.013$ $0.001$
+Ours $\bm{0.135}$ $\bm{0.959}$ $0.385$ $1.45$ $\bm{0.204}$ $\bm{0.056}$ $\bm{0.013}$ $0.002$
+IV Experiments
+IV-A Datasets and Metrics
+We trained and evaluated our model on both the KIT-ML [16] and HumanML3D [15] datasets.
+The KIT-ML dataset contains 3,911 motions and 6,363 natural language annotations, while HumanML3D is derived from and reannotates the HumanAct12 [69] and AMASS [70] datasets, comprising a total of 14,616 motion sequences and 44,970 descriptions.
+We follow the evaluation metrics used in previous motion generation works [24, 61, 64], including Frechet Inception Distance (FID), Precision, Diversity, Multimodality, and Multi-Modal Distance.
+Additionally, we consider physics-related metrics such as skating ratio, ground floating, ground penetration, and foot clipping, which have also been used in prior studies [22, 25, 23].
+Furthermore, we introduce two novel metrics derived from Fourier analysis: Velocity Consistency (VC) and Acceleration Consistency (AC).
+Our motivation stems from the observation that while the Smoothness metric captures global temporal coherence, it overlooks the inherent kinematic relationship between stable body parts and agile extremities in human motion.
+For a motion sequence represented as $\Phi\in\mathbb{R}^{T\times 3d}$, the metrics are formally defined as:
+VC $=\frac{1-\|\text{FFT}(\dot{\Phi})_{[:thre_{\text{low}}]}\|-\|\text{FFT}(\dot{\Phi})_{[thre_{\text{high}}:]}\|}{\|\text{FFT}(\dot{\Phi})_{[:thre_{\text{low}}]}\|+\|\text{FFT}(\dot{\Phi})_{[thre_{\text{high}}:]}\|},$ (12)
+AC $=\frac{1-\|\text{FFT}(\ddot{\Phi})_{[:thre_{\text{low}}]}\|-\|\text{FFT}(\ddot{\Phi})_{[thre_{\text{high}}:]}\|}{\|\text{FFT}(\ddot{\Phi})_{[:thre_{\text{low}}]}\|+\|\text{FFT}(\ddot{\Phi})_{[thre_{\text{high}}:]}\|}.$
+Spectral ratio analysis enables VC and AC to effectively characterize overall motion patterns by precisely quantifying the balance between high-frequency components (associated with rapid limb movements) and low-frequency components (correlated with stable torso dynamics).
+This frequency-domain approach provides a more accurate assessment of motion physicality compared to traditional smoothness metrics.
+Since the dataset includes motions like “climbing stairs” (a foot is floating from the ground) and “jumping while spinning” (abrupt acceleration), all these physics-based metrics should be made as consistent as possible with ground-truth instead of being blindly minimized to meet the characteristics of dataset.
+IV-B Evaluation Metrics
+In designing physical plausibility metrics, we draw on the Skating Ratio proposed in Gmd [26], the Floating, Ground Penetration, and Foot Clipping metrics from ReinDiffuse [22], and the Speed and Smoothness metrics introduced by Fischer [14].
+The specific definitions of these metrics are as follows:
+  • Skating Ratio: Foot skating refers to the situation where a person’s foot is in contact with the ground but still has velocity.
+Given a human motion $X$, the Foot Skating Ratio for the $i$-th frame can be expressed as:
+    $FS^{i}(X)=\exp\left(-||\dot{X^{i}_{ft}}\cdot f^{i}_{S}\cdot f^{i}_{V}\cdot f^{i}_{\bar{V}}||_{2}\right),$ (13)
+    where $\dot{X^{i}_{ft}}$, $f^{i}_{S}$, $f^{i}_{V}$, and $f^{i}_{\bar{V}}$ represent the foot velocity, the contact label, the instantaneous velocity threshold of the foot, and the average velocity threshold of the foot at the $i$-th frame, respectively.
+  • Floating: Floating refers to a scenario where the lowest joint of the generated motion is not in contact with the ground. Specifically, the Floating for the $i$-th frame can be represented as:
+    $Float^{i}(X)=\exp\left(-||(X^{i}_{h}-h_{ground})\cdot f^{i}_{F}||_{2}\right),$ (14)
+    where $X^{i}_{h}$ denotes the y-axis coordinate of the lowest point of the action at the $i$-th frame, $h_{ground}$ is a threshold hyperparameter, set to 5cm in practice, and $f^{i}_{F}$ is a flag that equals 1 when $X^{i}_{h}>h_{ground}$ and 0 otherwise.
+  • Ground Penetration: Ground Penetration refers to the situation where the lowest joint of the generated motion penetrates below the ground.
+The Ground Penetration for the $i$-th frame can be expressed as:
+    $GP^{i}(X)=\exp\left(-||(h_{ground}-X^{i}_{h})\cdot f^{i}_{P}||_{2}\right),$ (15)
+    where $f^{i}_{P}$ is a flag that equals 1 when $X^{i}_{h}<h_{ground}$ and 0 otherwise.
+  • Foot Clipping: Foot Clipping refers to a situation where the two feet are too close together, causing mesh penetration. The Foot Clipping for the $i$-th frame can be expressed as:
+    $FC^{i}(X)=\exp\left(-||(X^{i}_{lf}-X^{i}_{rf})\cdot f^{i}_{C}||_{2}\right),$ (16)
+    where $X^{i}_{lf}$ and $X^{i}_{rf}$ are the joint coordinates of the left and right feet, and $f^{i}_{C}$ is a threshold function.
+$f^{i}_{C}=1$ when the distance between the two feet exceeds a threshold, which is set to 5cm in practice.
+  • Speed: Speed can be used to calculate the amplitude of human motion. The calculation for Speed at the $i$-th frame is given by:
+    $Speed^{i}(X)=\frac{||X^{i}||_{2}}{\Delta_{t}},$ (17)
+    where $\Delta_{t}$ is the sampling frame rate of the dataset, which is 0.05 for HumanML3D and 0.08 for KIT-ML.
+  • Smoothness: Smoothness serves as a measure of fluctuations in motion acceleration.
+Models of inferior quality often exhibit low smoothness in their generated motions.
+The calculation for Smoothness at the $i$-th frame is given by:
+    $Smooth^{i}(X)=\frac{|\ddot{X^{i}}-\ddot{X^{i-1}}|}{\Delta_{t}}.$ (18)
+It is important to note that all the aforementioned metrics should ideally be as close as possible to the data set.
+Since the text-to-motion datasets include motions such as crawling, climbing stairs, and swimming, metrics like Skating Ratio and Floating are not necessarily better when smaller.
+For the KIT-ML dataset [16], Ground Penetration and Clipping are also not zero, which may be related to the quality of the KIT-ML dataset.
+IV-C Quantitative Results
+Figure 4: User Study Results.
+User Study.
+We conducted a user study in which participants rated four aspects of the generated motions on a scale of 1 to 5: “Generation Quality”, “Motion Smoothness”, “Physical Plausibility” and “Text Alignment”, using the HumanML3D test set.
+Thirty participants were invited to evaluate and compare the motion generation results of StableMoFusion + BioVAE, StableMoFusion, and the ground truth across 20 prompts.
+The results show that our model outperforms StableMoFusion in generation quality, motion smoothness, physical plausibility, and semantic alignment.
+In particular, our method is capable of generating results nearly indistinguishable from the ground truth in some cases, highlighting its effectiveness.
+Figure 5: User study interface for evaluating on HumanML3D. To ensure unbiased feedback, the samples were presented in random order to avoid potential scoring inertia among participants.
 Main Results.
-Tab.
-1 and Tab.
-2 compare our method with existing approaches across both physical plausibility metrics (e.g., smoothness, foot sliding) and traditional generation metrics (FID, Diversity).
-Our model achieves state-of-the-art results in biomechanical authenticity while simultaneously advancing traditional motion quality, outperforming prior works across multiple dimensions.
-Experiments reveal critical limitations in existing paradigms: GPT-style generative models tend to over-smooth motions, failing to capture realistic dynamics, while diffusion-based methods exhibit temporal discontinuity—likely due to their implicit treatment of spatiotemporal relationships without explicit temporal modeling.
-By integrating our physics-aware network into the diffusion framework, we address these issues: the network’s frame-wise acceleration recalculation enhances temporal dependency learning, directly improving physical metrics.
-Concurrently, the introduction of biomechanical priors reduces solution space complexity, enabling superior performance in traditional metrics without compromising motion diversity.
+Table I and Table II compare our framework with existing approaches in both physical plausibility metrics (e.g., smoothness, foot sliding) and traditional generation metrics (FID, Diversity).
+Our approach achieves state-of-the-art results in biomechanical authenticity while simultaneously advancing traditional motion quality, outperforming prior works across multiple dimensions.
+Experiments reveal critical limitations in existing paradigms: GPT-style generative models tend to over-smooth motions, failing to capture realistic dynamics, while diffusion-based methods exhibit temporal discontinuity, probably due to their implicit treatment of spatio-temporal relationships without explicit temporal modeling.
+In contrast, our method leverages the residual joint training of BioVAE and diffusion models, enabling the diffusion process to effectively capture acceleration information during training.
+Meanwhile, the incorporation of biomechanical priors reduces the complexity of the solution space, allowing our model to achieve superior performance on traditional metrics without compromising motion diversity.
 Figure 6: Qualitative comparison of the state-of-the-art methods.
-Quantitative evaluations demonstrate that BioMoDiffuse achieves superior language-motion alignment, enhanced temporal smoothness, and improved physical plausibility compared to current approaches.
-4.3 Ablation Study
-Table 3: Ablation Study on traditional metrics.
-we evaluate the impact of adding $L_{accel}$ loss, incorporating the Physical Network, and introducing Physics Guidance on the performance of our method.
-All experiments were repeated 20 times with different random seeds, and the mean results are reported in the tables.
+The results demonstrate that our method achieves superior language-motion alignment, enhanced temporal smoothness, and improved physical plausibility compared to current approaches.
+IV-D Ablation Study
+TABLE III: Ablation study on traditional metrics. All experiments were repeated 20 times with different random seeds, and the mean results are reported in the tables.
 […]
@@
 Ours $\times$ $\times$ $\times$ $0.090$ $2.824$ $9.590$ $1.849$ $0.542$ $0.739$ $0.832$
-$\checkmark$ $\times$ $\times$ $0.129$ $2.835$ $\bm{9.659}$ $1.882$ $0.537$ $0.734$ $0.827$
-$\checkmark$ $\checkmark$ $\times$ $\bm{0.066}$ $\bm{2.783}$ $9.596$ $\bm{1.948}$ $\bm{0.549}$ $\bm{0.746}$ $\bm{0.836}$
+$\checkmark$ $\times$ $\times$ $0.085$ $\bm{2.760}$ $9.764$ $1.835$ $\bm{0.551}$ $\bm{0.747}$ $\bm{0.839}$
+$\checkmark$ $\checkmark$ $\times$ $\bm{0.066}$ $2.783$ $9.596$ $\bm{1.948}$ $0.549$ $0.746$ $0.836$
+$\checkmark$ $\times$ $\checkmark$ $0.077$ $2.803$ $9.752$ $1.873$ $0.545$ $0.740$ $0.832$
 $\checkmark$ $\checkmark$ $\checkmark$ $0.071$ $2.784$ $9.567$ $1.919$ $0.547$ $0.743$ $0.835$
-Table 4: Ablation Study on physical plausible metrics.
+TABLE IV: Ablation study on physical plausible metrics.
 […]
@@
 Ours $\times$ $\times$ $\times$ $0.292$ $2.220$ $0.448$ $0.74$ $0.234$ $0.074$ $0.043$ $0.002$
-$\checkmark$ $\times$ $\times$ $0.144$ $1.022$ $0.437$ $1.24$ $0.228$ $0.076$ $0.052$ $0.003$
-$\checkmark$ $\checkmark$ $\times$ $0.135$ $\bm{0.954}$ $0.421$ $1.39$ $0.227$ $0.060$ $0.026$ $0.002$
-$\checkmark$ $\checkmark$ $\checkmark$ $\bm{0.135}$ $0.959$ $\bm{0.385}$ $\bm{1.45}$ $\bm{0.204}$ $\bm{0.056}$ $\bm{0.013}$ $\bm{0.002}$
-Our method introduces three key innovations to the diffusion framework: acceleration loss, a physics network grounded in Euler-Lagrange equations, and physics-guided sampling.
-Ablation studies under identical training settings Tab.
-3 and Tab.
-4 validate their contributions.
-While the acceleration loss enhances kinematic feature extraction by expanding the motion representation dimension from 263 to 329, it initially degrades generation quality (e.g., FID deteriorates by 43.3%) due to increased solution space complexity.
-The proposed physics network addresses this by embedding biomechanical constraints, reformulating motion generation as a constrained optimization problem.
-This reduces spatial complexity while enforcing rigid-body dynamics, leading to a 48.8% FID improvement with only a 6.68% parameter increase over the UNet baseline.
-Crucially, the network’s Euler-Lagrange foundation enables simultaneous gains in physical plausibility and generation quality.
-Finally, physics-guided sampling prioritizes biomechanical fidelity during denoising, stabilizing acceleration fitting and further improving physical metrics, albeit with a slight trade-off in generation scores—a phenomenon consistent with CFG principles.
-4.4 Motion Refining Results
-Table 5: Comparative results for pose refinement.
-We trained and tested the WHAM [60] baseline and WHAM $+$ our method using the official code on the 3DPW [70] and RICH [21] datasets.
+$\checkmark$ $\times$ $\times$ $0.161$ $1.205$ $0.418$ $1.23$ $0.227$ $0.056$ $0.026$ $0.002$
+$\checkmark$ $\checkmark$ $\times$ $0.135$ $0.954$ $0.421$ $1.39$ $0.227$ $0.060$ $0.026$ $0.002$
+$\checkmark$ $\times$ $\checkmark$ $\bm{0.134}$ $\bm{0.944}$ $0.425$ $1.33$ $0.211$ $0.056$ $0.039$ $0.002$
+$\checkmark$ $\checkmark$ $\checkmark$ $0.135$ $0.959$ $0.385$ $\bm{1.45}$ $\bm{0.204}$ $\bm{0.056}$ $\bm{0.013}$ $\bm{0.002}$
+Our proposed BioVAE not only incorporates additional biomechanical priors but also achieves stable joint training by modifying the diffusion loss function.
+Ablation studies in identical training settings (Table III and Table IV) demonstrate the critical contributions of the BioVAE architecture, the biomechanical priors, and the newly introduced loss term.
+Specifically, BioVAE substantially improves motion generation performance on traditional metrics, while the biomechanical priors and BioVAE loss each independently enhance the physical plausibility of the generated motions.
+When combined, these components exhibit even greater improvements, highlighting the effectiveness and advancement of our framework.
+Figure 7: Additional samples from our text-to-motion synthesis approach, generated with text prompts from the HumanML3D test set.
+The results demonstrate that our approach effectively understands diverse textual semantics and generates physically plausible motions.
+IV-E Motion Estimation and Editing
+TABLE V: Comparative results for pose estimation.
+We used WHAM [4] and various diffusion models as the base framework.
+Building on the official WHAM implementation, we evaluated the impact of incorporating the residual-based BioVAE on pose estimation performance using the 3DPW [17] and RICH [18] datasets.
 Bold numbers denote the most accurate method in each column.
 Accel is in $m/s^{2}$, all other errors are in $mm$.
+Method 3DPW RICH
 […]
-3DPW WHAM $36.34$ $61.46$ $70.64$ $6.52$
-+ Inversion $\bm{36.13}$ $\bm{61.11}$ $\bm{70.52}$ $\bm{5.55}$
-RICH WHAM $50.67$ $83.00$ $95.32$ $5.35$
-+ Inversion $\bm{50.60}$ $\bm{81.80}$ $\bm{94.84}$ $\bm{4.06}$
-As shown in Tab.
-5, we validate our model’s motion refinement capability using WHAM [60], a state-of-the-art pose estimation framework that leverages temporal modeling to achieve leading results on metrics like acceleration error.
-By integrating our physics-aware network with WHAM via diffusion inversion, we significantly improve motion smoothness while also advancing pose prediction accuracy to new state-of-the-art levels.
-This indicates that our model, equipped with diffusion inversion, can be applied after any pose estimation module to refine the motion sequence, resulting in more fluent and physically plausible results.
-These results demonstrate that our physical network effectively enhances temporal continuity and overall robustness, even without acceleration loss supervision.
-This dual capability underscores the versatility of our approach, showcasing its ability to refine both generated and estimated motions by enforcing physical plausibility at the intersection of kinematics and dynamics.
-4.5 Ablation Experiments on Speed Control Signal
-Table 6: Ablation Study on weight of Speed Control Signal.
-We evaluate the impact of different Speed Control Signal on generation quality and physical plausibility metrics.
-All experiments use the HumanML3D dataset.
+WHAM 36.34 61.46 70.64 6.52 50.67 83.00 95.32 5.35
+MDM Inversion 36.48 62.06 71.36 6.41 51.76 83.62 96.53 5.28
+StableMoFusion Inversion $\bm{36.06}$ 61.85 71.26 6.29 51.64 83.12 96.53 5.11
+StableMoFusion + BioVAE Inversion 36.13 $\bm{61.11}$ $\bm{70.52}$ $\bm{5.55}$ $\bm{50.60}$ $\bm{81.80}$ $\bm{94.84}$ $\bm{4.06}$
+As shown in Table V, we evaluate the capability of our framework for motion refinement and pose estimation using WHAM [4], a state-of-the-art method in the pose estimation domain.
+WHAM incorporates sequential modeling and temporal information, achieving leading results in acceleration-related metrics.
+Our results demonstrate that integrating BioVAE with diffusion models and pose estimation frameworks can significantly improve the smoothness of human motion while also enhancing pose estimation accuracy, achieving state-of-the-art performance across all evaluated metrics.
+These experiments indicate that our physics-aware network not only reduces the complexity of the solution space and enhances the physical plausibility but also substantially improves the smoothness of the motion and the overall performance even without relying on acceleration loss supervision.
+This highlights the versatility of our approach to simultaneously refining human motion and advancing pose estimation results.
+TABLE VI: Motion Editing via Acceleration Modification.
+We manually edited the acceleration values in BioVAE to evaluate the effects of increasing or decreasing acceleration on the generated motions.
+All experiments were conducted on the HumanML3D dataset.
 […]
@@
 $2.0$ $0.124$ $2.869$ $0.824$ $0.136$ $0.942$ $0.447$ $1.26$ $0.321$ $0.045$ $0.013$ $0.002$
-Our physical network incorporates a speed control signal within the autoregressive block.
-Experiments reveal that applying this signal only during inference (via input weighting) dynamically scales motion amplitude, with stronger signals producing larger movements.
-However, three critical observations emerge: (1) Traditional generation metrics (e.g., FID) remain unaffected by signal strength, aligning with the convention that model parameters require consistency between training and testing; (2) Certain physical metrics, such as ground penetration, foot clipping, and velocity/acceleration consistency, also show minimal sensitivity; (3) Metrics tied to acceleration magnitude—smoothness and skating ratio—decrease with stronger signals, while ground floating increases.
-We attribute this to the explicit acceleration modeling in our framework: since the network learns absolute acceleration values (rather than relative relationships), scaling the input signal alters acceleration magnitudes, thereby directly impacting physics-based metrics.
-This demonstrates our model’s unique ability to modulate motion amplitude through acceleration control while preserving generation quality—a trade-off avoided by conventional methods.
-Furthermore, these findings guide targeted optimization of physical plausibility by tuning acceleration-related parameters.
-5 Conclusion
-In this paper, we present BioMoDiffuse, a motion generation model built upon a physical network derived from extended Euler-Lagrange equations.
-This model incorporates the physical network into pose estimation using DDIM Inversion.
-BioMoDiffuse combines physics guidance with a state-of-the-art UNet backbone architecture, while simultaneously learning from electromyographic (EMG) signals and acceleration data.
-Experimental results indicate that BioMoDiffuse surpasses previous models in both generation quality and physical plausibility.
-However, due to current dataset limitations, BioMoDiffuse has not yet been tested on higher-frame-rate datasets with more comprehensive annotations.
-We leave it for future exploration.
-Supplementary Material
-Video:
-We have provided supplemental videos in our supplemental material.
-In the supplementary video, we demonstrate the following: 1) the differences between the human skeletons generated by our method and those by StableMoFusion [22], 2) the differences in the skinned human models generated by our method and StableMoFusion, and 3) the text prompts used to generate these motions.
-All generations were produced using the same code under identical random seed conditions (seed=0).
-We reconstructed the meshes following the rendering approach provided by MLD [5].
-Code:
-The code has also been included in the supplementary materials, and detailed execution instructions can be found in the accompanying README file.
-Appendix A Evaluation Metrics
-In the design of physical plausibility metrics, we refer to the design of the Skating Ratio in Gmd [29], as well as the designs for Floating, Ground Penetration, and Foot Clipping in ReinDiffuse [18], as well as the Speed and Smoothness metrics provided by Fischer [10].
-The specific definitions of these metrics are as follows:
-  • Skating Ratio: Foot skating refers to the situation where a person’s foot is in contact with the ground but still has velocity.
-Given a human motion $\Theta$, the Foot Skating Ratio for the $i$-th frame can be expressed as:
-    $FS^{i}(\Theta)=\exp\left(-||\dot{\Theta^{i}_{ft}}\cdot f^{i}_{S}\cdot f^{i}_{V}\cdot f^{i}_{\bar{V}}||_{2}\right)$ (14)
-    where $\dot{\Theta^{i}_{ft}}$, $f^{i}_{S}$, $f^{i}_{V}$, and $f^{i}_{\bar{V}}$ represent the foot velocity, the contact label, the instantaneous velocity threshold of the foot, and the average velocity threshold of the foot at the $i$-th frame, respectively.
-  • Floating: Floating refers to a scenario where the lowest joint of the generated motion is not in contact with the ground. Specifically, the Floating for the $i$-th frame can be represented as:
-    $Float^{i}(\Theta)=\exp\left(-||(\Theta^{i}_{h}-h_{ground})\cdot f^{i}_{F}||_{2}\right)$ (15)
-    where $\Theta^{i}_{h}$ denotes the y-axis coordinate of the lowest point of the action at the $i$-th frame, $h_{ground}$ is a threshold hyperparameter, set to 5cm in practice, and $f^{i}_{F}$ is a flag that equals 1 when $\Theta^{i}_{h}>h_{ground}$ and 0 otherwise.
-  • Ground Penetration: Ground Penetration refers to the situation where the lowest joint of the generated motion penetrates below the ground.
-The Ground Penetration for the $i$-th frame can be expressed as:
-    $GP^{i}(\Theta)=\exp\left(-||(h_{ground}-\Theta^{i}_{h})\cdot f^{i}_{P}||_{2}\right)$ (16)
-    where $f^{i}_{P}$ is a flag that equals 1 when $\Theta^{i}_{h}<h_{ground}$ and 0 otherwise.
-  • Foot Clipping: Foot Clipping refers to a situation where the two feet are too close together, causing mesh penetration. The Foot Clipping for the $i$-th frame can be expressed as:
-    $FC^{i}(\Theta)=\exp\left(-||(\Theta^{i}_{lf}-\Theta^{i}_{rf})\cdot f^{i}_{C}||_{2}\right)$ (17)
-    where $\Theta^{i}_{lf}$ and $\Theta^{i}_{rf}$ are the joint coordinates of the left and right feet, and $f^{i}_{C}$ is a threshold function.
-$f^{i}_{C}=1$ when the distance between the two feet exceeds a threshold, which is set to 5cm in practice.
-  • Speed: Speed can be used to calculate the amplitude of human motion. The calculation for Speed at the $i$-th frame is given by:
-    $Speed^{i}(\Theta)=\frac{||\Theta^{i}||_{2}}{\Delta_{t}}$ (18)
-    where $\Delta_{t}$ is the sampling frame rate of the dataset, which is 0.05 for HumanML3D and 0.08 for KIT-ML.
-  • Smoothness: Smoothness can be used to measure the jittering of the model’s acceleration.
-Models with poor quality tend to exhibit poor Smoothness.
-The calculation for Smoothness at the $i$-th frame is given by:
-    $Smooth^{i}(\Theta)=\frac{|\ddot{\Theta^{i}}-\ddot{\Theta^{i-1}}|}{\Delta_{t}}$ (19)
-It is important to note that all of the aforementioned metrics should ideally be as close as possible to the dataset.
-Since the dataset includes motions such as crawling, climbing stairs, and freestyle swimming, metrics like Skating Ratio and Floating are not necessarily better when smaller.
-For the KIT-ML dataset [51], Ground Penetration and Clipping are also not zero, which may be related to the quality of the KIT-ML dataset.
-Appendix B Implementation Details
-For training, we employ DDPM [20] with 1,000 denoising steps ($T=1,000$) and linearly varying variances $\beta_{t}$ from 0.0001 to 0.02 in the forward process.
-The model is trained using AdamW [40] with an initial learning rate of 0.0002 and a weight decay of 0.01 for 50,000 iterations during pre-training, followed by 10,000 iterations of fine-tuning with a batch size of 64 on an NVIDIA A40 GPU.
-The learning rate is reduced by a factor of 0.9 every 5,000 steps.
-Gradient norms are clipped to a maximum value of 1 during gradient descent.
-For class-conditional generation (CFG) [19], text conditioning $c$ is set to $\emptyset$ for 10% of the samples, and the physical network is ignored with a probability of $P_{c}=10\%$.
-Additionally, we apply Exponential Moving Average (EMA) to smooth the parameter updates and reduce mutations and oscillations.
-For inference, we utilize the SDE variant of the second-order DPM-Solver++ [41] with Karras sigmas [28] for 10-step sampling. The scale for text conditioning CFG is set to 2.5.
-In our implementation of pose refinement, we used WHAM [60] as the baseline.
-However, the official WHAM code repository only provides versions trained on five datasets: AMASS [44], InstaVariety [27], MPI-INF-3DHP [45], Human3.6M [23], and 3DPW [70].
-It does not include a version that incorporates the BEDLAM [4] dataset for training.
-As a result, the baseline results in our work differ from those presented in the WHAM paper.
-Appendix C Network Architecture
-Table 7: Model Architecture of BioMoDiffuse on dataset HumanML3D.
-Model Module Details
-BioMo Diffuse Physical Network (1-1)
-• GRU layer
-• AccelNet:
-   - Motion MLP
-   - Muscle MLP
-   - Physics Decoder
-   - Muscle Decoder
-   - M_Net
-   - F_Net
-   - C_Net
-   - Pose2EMG
-Cond Unet1D (1-2)
-• Time MLP
-• Downsampling Blocks (×4) :
-   - Conv1D
-   - Cross Attention
-   - Residual Blocks
-• Mid Blocks (×2) :
-• Upsampling Blocks (×4) :
-   - Transposed Conv1D
-   - Residual Blocks
-Tab.
-7 illustrates the network architecture of BioMoDiffuse trained on HumanML3D [14].
-Compared to the baseline UNet, the addition of our physical network only increases the parameter count by $6.68\%$.
-However, this results in a significant improvement in the FID score, which decreases by $48.8\%$ (from 0.129 to 0.068).
-Other metrics also show notable improvements, demonstrating that the proposed physical network is both lightweight and effective.
-Appendix D Derivation of Euler-Lagrange equation
-The classical Euler-Lagrange equation is formulated as:
-$\frac{d}{dt}\left(\frac{\partial L}{\partial\dot{q}}\right)-\frac{\partial L}{\partial q}=Q,$ (20)
-where $L=T-V$ is the Lagrangian, $T$ denotes kinetic energy, $V$ represents potential energy, $q$ are generalized coordinates, and $Q$ encompasses non-conservative forces such as joint torques and external contact forces.
-By modeling the human body as a rigid-body system, its kinetic energy can be expressed as:
-$T=\frac{1}{2}\dot{q}^{T}M(q)\dot{q},$ (21)
-where $M(q)$ is the inertia matrix. The potential energy $V(q)$ is defined by the gravitational field. Substituting $L=T-V$ into the Euler-Lagrange equation yields:
-$\frac{d}{dt}\left(\frac{\partial T}{\partial\dot{q}}\right)-\frac{\partial T}{\partial q}+\frac{\partial V}{\partial q}=Q.$ (22)
-Substituting the explicit form of $T$, we derive:
-$M(q)\ddot{q}+\dot{M}(q)\dot{q}-\frac{1}{2}\dot{q}^{T}\frac{\partial M}{\partial q}\dot{q}+G(q)=Q,$ (23)
-where $G(q)=\frac{\partial V}{\partial q}$. Leveraging the symmetry of $M(q)$, we reorganize indices and employ Christoffel symbols to define the Coriolis matrix:
-$C(q,\dot{q})\dot{q}=\sum_{i,j}\frac{1}{2}\left(\frac{\partial M_{kj}}{\partial q_{i}}+\frac{\partial M_{ki}}{\partial q_{j}}-\frac{\partial M_{ij}}{\partial q_{k}}\right)\dot{q}_{i}\dot{q}_{j}.$ (24)
-It can be rigorously shown that $C(q,\dot{q})\dot{q}=\dot{M}(q)\dot{q}-\frac{1}{2}\dot{q}^{T}\frac{\partial M}{\partial q}$, demonstrating that the Coriolis term arises from the derivative of the inertia matrix and quadratic velocity coupling, which depend solely on the body’s pose.
-Figure 6: More samples from our BioMoDiffuse for text-to-motion synthesis. Samples generated with text prompts of HumanML3D’s test set. We recommend the supplemental video to see more motion results.
-Appendix E Comparison on Samplers
-Table 8: Comparative Evaluation of Motion Generation Methods. We compare different methods on generation quality, efficiency, and physical plausibility metrics using the HumanML3D dataset.
-[…]
-GT - - $0.002$ $2.974$ $0.797$ $0.124$ $0.682$ $0.358$ $2.60$ $0.205$ $0.057$ $0.000$ $0.000$
-DEIS $10$ $3.659$ $0.148$ $2.732$ $0.843$ $0.298$ $2.388$ $0.380$ $0.91$ $0.195$ $0.054$ $0.013$ $0.002$
-DEIS w/o Karras $10$ $3.615$ $0.146$ $2.723$ $0.844$ $0.271$ $2.285$ $0.381$ $0.88$ $0.195$ $0.051$ $0.013$ $0.002$
-DPMSolver++ w/o Karras $10$ $3.664$ $0.126$ $2.755$ $0.840$ $0.147$ $1.071$ $0.374$ $1.39$ $0.196$ $0.050$ $0.012$ $0.001$
-DPMSolver++ (20) $20$ $7.610$ $0.090$ $2.781$ $0.834$ $0.143$ $0.969$ $0.382$ $1.44$ $0.203$ $0.059$ $0.013$ $0.003$
-DPMSolver++ (10) $10$ $2.553$ $0.071$ $2.784$ $0.835$ $0.135$ $0.959$ $0.385$ $1.45$ $0.204$ $0.056$ $0.013$ $0.002$
-DDIM $50$ $19.82$ $0.123$ $2.741$ $0.839$ $0.132$ $0.933$ $0.368$ $1.57$ $0.185$ $0.054$ $0.012$ $0.002$
-DDPM $1000$ $290.30$ $0.088$ $2.771$ $0.833$ $0.147$ $0.981$ $0.385$ $1.43$ $0.209$ $0.060$ $0.013$ $0.003$
-In this section, we present the impact of different samplers on the generation performance of BioMoDiffuse.
-We evaluated the following samplers on our model: DDPM [20] , DDIM [62], DPMSolver [41], and DEIS [84].
-Additionally, we incorporated Karras Sigma into DEIS and DPMSolver.
-Karras Sigma corresponds to constant-velocity thermal diffusion, which has been shown to ensure faster and higher-quality generation in the field of image synthesis.
-As shown in Tab.
-8, we observe that DEIS fails to generate sufficiently smooth motions, while among the other samplers, DPMSolver achieves superior results in terms of FID and physical plausibility metrics.
-Furthermore, the inclusion of Karras Sigma contributes to improved FID scores.
-Based on these findings, we ultimately selected DPMSolver with Karras Sigma as the preferred configuration for our framework.
-Appendix F User Study
-Figure 7: User study interface for evaluating StableMoFusion and our model on HumanML3D.
-In the user study, we generated multiple samples using the evaluation code on the HumanML3D test set, filtering out semantically redundant examples to create a final set of 20 distinct samples.
-Participants were asked to rate the generated motions on a 1–5 scale across four dimensions: ”Generation Quality,” ”Motion Smoothness,” ”Physical Plausibility,” and ”Text Alignment.” We invited 30 users to compare the motion generation results of our model, StableMoFusion, and Ground Truth based on these 20 prompts.
-To ensure unbiased feedback, the samples were presented in random order, preventing participants from identifying which motions were generated by our model.
-Appendix G More Qualitative Results
-Fig.
-6 demonstrates the generation performance of our model across a broader set of prompts, while Fig.
-8 provides a comparative analysis of our model against state-of-the-art methods on an extended set of prompts.
-The results indicate that our model consistently achieves superior performance across a wider range of motion types.
-Figure 8: The comparison of the state-of-the-art methods on text-guided motion synthesis task.
-Samples generated with text prompts of HumanML3D’s test set.
-We recommend the supplemental video to see more motion results.
+By modifying the acceleration values in BioVAE for ablation studies, we observed the following findings: (1) Traditional generation metrics generally do not improve with changes in acceleration magnitude, which explains why generative models require parameter consistency between training and testing; (2) Certain physical metrics, such as ground penetration, foot clipping, and velocity/acceleration consistency, also show minimal sensitivity to acceleration changes; (3) Some physical metrics exhibit a strong correlation with acceleration magnitude: smoothness and skating ratio decrease as acceleration increases, while ground floating increases with higher acceleration values.
+We attribute these phenomena to the explicit representation of acceleration in our model.
+When we modify the absolute magnitude of the input acceleration, the model shows little impact on traditional generation metrics.
+However, because our method explicitly models acceleration, altering its absolute magnitude affects the generated motion amplitude and, in turn, influences acceleration-related physical plausibility metrics.
+This highlights the advantage of our approach: the ability to regulate acceleration to control the physical properties of generated motions.
+In addition, these insights provide guidance for optimizing specific physical metrics by tuning acceleration-related parameters.
+IV-F Implementation Details
+All experiments were conducted on a single NVIDIA A40 GPU.
+For our implementation of motion generation, we adopt StableMoFusion as the baseline.
+We employ DDPM [39] with 1,000 denoising steps ($T=1{,}000$) and linearly varying variances $\beta_{t}$ from 0.0001 to 0.02 in the forward process.
+The model is trained using AdamW [71] with an initial learning rate of 0.0002 and a weight decay of 0.01 for 50,000 iterations during pre-training, followed by 10,000 iterations of fine-tuning with a batch size of 64.
+Additionally, we apply an Exponential Moving Average (EMA) to smooth parameter updates and reduce instability and oscillations.
+During inference, we utilize the SDE variant of the second-order DPM-Solver++ [72] with Karras sigmas [73] for 10-step sampling.
+We set the CFG scale for text conditioning to 2.5.
+For our implementation of pose refinement, we adopted WHAM [4] as the baseline.
+However, the official WHAM code repository provides models trained only on five datasets: AMASS [70], InstaVariety [74], MPI-INF-3DHP [75], Human3.6M [76], and 3DPW [17], and does not include a version trained with the BEDLAM [77] dataset.
+Consequently, the baseline results reported in our work differ slightly from those presented in the original WHAM paper.
+Figure 8: Editing generated human motions to different movement amplitudes.
+By manually modifying the acceleration values in BioVAE, controllable motion generation can be achieved.
+Adjusting the acceleration for a single joint or all joints allows the generated motions to exhibit varying amplitudes.
+V Conclusion
+In this paper, we propose BioVAE, a biomechanics-aware framework designed for a wide range of tasks, including motion generation, pose estimation, motion refinement, and motion editing.
+BioVAE is inspired by extended Euler-Lagrange equations and achieves deep integration with the diffusion process through a carefully designed loss function.
+Furthermore, it incorporates joint learning from electromyographic (EMG) signals and acceleration data.
+Experimental results demonstrate that our approach outperforms previous models in both generation quality and physical plausibility.
+However, data limitations have so far prevented us from evaluating our approach on high-frame-rate datasets with richer annotations.
+Addressing this limitation in future work could enable BioVAE to deliver even stronger results for practical industrial applications.